### v2

# Observability-Aware Control for Quadrotor Formation Flight with Range-only Measurement

Abstract
Cooperative Localization (CL) is a promising approach to achieve safe quadrotor formation flight through precise positioning via low-cost inter-drone sensors. This paper develops an observability-aware control principle tailored to quadrotor formation flight with range-only inter-drone measurement. The control principle is based on a novel approximation of the local observability Gramian (LOG), which we name the Short-Term Local Observability Gramian (STLOG). The validity of STLOG is established by a proof of the link between local observability and estimation precision. We propose the Observability Predictive Controller (OPC), an implementation of our control principle under a receding-horizon framework, which optimizes a metric of the STLOG to maximize the minimum precision improvement along a trajectory. Monte Carlo simulations and experimental flight tests are conducted on a pair of quadrotors performing formation flight. The results show that the OPC improves positioning precision and estimator confidence, confirming the practical utility of the proposed approach.
Nomenclature
$d\in\mathbb{R}^{+}$  = The distance between the leader and the follower, measured by the follower
DfDf  = The derivative of a unary function $f$ with respect to its argument
$E[\mathbf{x}],\mathbb{V}[\mathbf{x}]$  = The expected value and variance of a random variable $\mathbf{x}$
$f^{1},f^{2}\in\mathbb{R}^{+}$  = The thrust-per-unit-mass delivered along each quadrotor’s body z-axis
$\mathbf{H}$  = The observation matrix, which is the Jacobian of the observation model
$\mathcal{I}^{\star}(\cdot):\mathbb{R}^{3\times 1}\to\mathbb{H}$  = An operator transforming a three-dimensional vector into a quaternion with zero real part
$L^{(r)}_{\mathbf{f}}g$  = The $r$th Lie derivative of a function $g$ along a vector field $\mathbf{f}$
$\mathbf{K}(t,t^{\prime}),\ \mathcal{K}(\cdot)$  = The covariance function of the sensor noise process and its associated integral operator
$\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$  = A multivariate normal distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$
$p(\mathbf{x})$  = The probability density function of a random variable $\mathbf{x}$
$\check{\mathbf{P}}_{k},\hat{\mathbf{P}}_{k}$  = Prior and posterior error covariance matrix at timestep $k$
$\mathbf{q}\in\mathcal{S}^{3}$  = The quaternion of the follower relative to the leader
$\mathbf{r}\in\mathbb{R}^{3\times 1}$  = The position of the leader relative to the follower, expressed in the follower’s frame
$\mathbf{R}(\cdot):\mathcal{S}^{3}\to\mathbb{R}^{3\times 3}$  = The rotation matrix corresponding to a unit quaternion
$T$  = The time horizon over which observability is evaluated
$T_{k},T_{k+1}$  = The kth timestep under the OPC discretization scheme
$\boldsymbol{\Sigma}$  = Sensor noise covariance matrix, assumed to be time-invariant
$\mathbf{v}\in\mathbb{R}^{3\times 1}$  = The velocity of the leader relative to the follower, expressed in the follower’s frame
$\delta t$  = The correlation timescale of the observation noise
$\Delta T$  = Discretization step size for the system dynamics and predictive control
$\lambda_{\min}(\cdot),\sigma_{\min}(\cdot)$  = The minimum Eigenvalue and Singular value of matrices
$\phi_{t}$  = The flow of a dynamical system up to time $t$
$\boldsymbol{\Phi}_{t_{0},t_{1}}$  = The state transition matrix evaluated from $t_{0}$ up to time $t_{1}$
$\boldsymbol{\eta}(t)$  = A Gaussian Process representing observation noise
$\boldsymbol{\omega}^{l},\boldsymbol{\omega}^{f}\in\mathbb{R}^{3\times 1}$  = The body rate of each quadrotor, expressed in each vehicle’s own frame
$\mathbf{1},\mathbf{1}_{i}$  = The identity matrix and its $i$th column (c. f. the $i$th standard basis vector)
$\mathbf{0}_{m\times n}$  = A $m$-times-$n$ matrix filled with zeroes
$\times:\mathbb{R}^{3\times 1}\to\mathbb{R}^{3\times 3}$  = An operator trasforming a three-dimensional vector into a skew-symmetric matrix
$\otimes:\mathbb{H}\times\mathbb{H}\to\mathbb{H}$  = The quaternion product operator
1 Introduction
Cooperative localization (CL) or cooperative location system (CLS) allows multiple agents to work together to determine their positions more accurately than they could individually. It is one promising approach to support autonomous multi-quadrotor systems performing formation flight for applications such as search and rescue, surveillance, and environmental monitoring. CL uses inter-drone sensors such as ultra-wideband (UWB) range systems [1] to share information, improving the precision of positioning over short distances between drones in tight formations, or serves as a low-cost alternative to Global Navigation Satellite Systems (GNSS) when it becomes unreliable in dynamic and complex environments. However, a specific challenge arises in the range-only CLS of the leader-follower configuration: the follower drone(s) cannot determine its spatial position relative to the leader from a one-dimensional range measurement. Therefore, cooperative motion strategies must be carefully redesigned to ensure that the follower drone can determine its full state from range observations, preventing unbounded growth of localization uncertainty as depicted in Fig. 1.
Observability-aware trajectory optimization addresses this challenge by generating motion patterns to maximize the information gained from motion-coupled observations for precise state reconstruction. This approach has been proven to contribute to reducing the covariance of EKF estimation [2] — a classic problem in CL [3, 4, 5] that was investigated in our previous work [6]. The key to quantifying observability lies in the metrics of the LOG (local observability Gramian), which can be approximated numerically by the empirical local observability Gramian (ELOG) [7], or through the Lie derivatives of the observation model in the format of the expanded empirical local observability Gramian (E2LOG) [8, 9]. The latter approach enables observability-aware optimization to solve a critical issue with optimizing first-order metrics like the EKF covariance matrix, the Fisher information matrix (FIM), and the Cramér-Rao Lower Bound (CRLB): if parts of the system state are indirectly observed, their corresponding columns in the observation matrix will be constantly zero. This leads to rank-deficiency in the FIM, a part of the EKF covariance remaining unchanged from the initial prior covariance, and an undefined CRLB — a failure to quantify information about indirectly observed states in every case. A mitigation strategy — integrating the FIM over a trajectory — is still less effective than using the LOG to quantify trajectory-dependent information, since the time-integrated FIM only approaches the ELOG under ideal limits [10, 11]. The limitations of alternatives justify using the LOG to quantify the informativeness of motion-coupled observations, particularly in systems with indirect measurements, such as range-only quadrotor CLS.
[…]
(a) The follower enjoys improved positioning by using an observability-aware control strategy
[…]
(b) The follower suffers ambiguous positioning when following a uniform trajectory
Figure 1: Example configurations of a CL-aided leader-follower quadrotor team
Despite the theoretical maturity of the LOG and the development of approximating frameworks, applying observability-aware optimization to quadrotor CLS remains challenging. This is primarily due to the reliance on mechanically simplified models in existing work. Such models may distort the LOG metric, misrepresent couplings between motion and observability, and result in control strategies that are dynamically unfeasible for real quadrotors. In range-based relative localization, Salaris [12] applied observability-aware optimization based on metrics of the LOG to range-only landmark-based localization. The constructibility Gramian (CG) was later proposed [13] to improve information quantification, and De Carli [14] integrated it into a perception-aware path planner to improve multi-agent positioning. Lastly, observability-based optimization has been extended beyond offline trajectory design into the area of real-time optimal control [15, 16, 17]. However, these approaches rely on simplified dynamics, including planar quadrotors [13], first-order kinematic systems [14], or unicycles [15, 16, 17], all of which exhibit deficiencies in the actual dynamics of the quadrotor. For instance, unicycle models assume direct control over translational velocity aligned with heading, whereas quadrotor translation is coupled with roll and pitch (but not yaw). This mismatch may degrade the fidelity of observability metrics and lead to trajectories that are either suboptimal or infeasible, undermining the effectiveness of observability-aware optimization. Directly introducing full quadrotor dynamics when approximating the LOG can resolve this modeling gap, but it introduces significant computational demands that hinder real-time planning and control. These limitations motivate our development of a novel observability-aware control framework for range-only quadrotor CLS designed to ensure feasibility, effectiveness, and computational efficiency.
This paper presents an observability-aware control framework for quadrotor formation flight using range-only CL. We introduce the concept of short-term local observability Gramian (STLOG), a novel point-wise approximation of the LOG that enables real-time optimization. We then formulate an Observability-aware control principle that maximizes STLOG-based observability over the trajectory horizon. Our main contributions address the challenge of performing real-time observability-aware control under realistic quadrotor dynamics and dimensionally-limited range-only sensing:
  1. 1.
    We derive the STLOG and incorporate it into an observability-aware optimal control principle, thereby refining state estimation along the least observable directions.
  2. 2.
    We apply our controller to quadrotors in formation flight, demonstrating that observability-optimal trajectories improve positioning performance in Monte-Carlo simulations with realistic noise models, and flight experiments.
The remainder of the paper is organized as follows. In Section 2, we describe the range-only quadrotor CLS and the trajectory optimization problem. In Section 3, we study our CLS through classical observability analysis. We introduce the STLOG in Section 4, then incorporate it into our control principle in Section 5. Section 6 presents Monte-Carlo simulation and flight experimental results. Finally, Section 7 concludes the paper and discusses future work.
2 Problem Statement
2.1 Mathematical Preliminaries
[…]
$\mathbf{a}^{\times}=\begin{bmatrix}0&-a_{3}&a_{2}\\ a_{3}&0&-a_{1}\\ -a_{2}&a_{1}&0\end{bmatrix},\quad\mathbf{q}\otimes\mathbf{p}=\begin{bmatrix}q_{4}\mathbf{p}+\mathbf{q}_{1:3}^{\times}&\mathbf{q}_{1:3}\\ -\mathbf{q}_{1:3}&q_{4}\end{bmatrix}\begin{bmatrix}\mathbf{p}_{1:3}\\ p_{4}\end{bmatrix},\quad\mathbf{R}(\mathbf{q})=\mathbf{1}+2q_{4}\mathbf{q}_{1:3}^{\times}+2{\mathbf{q}_{1:3}^{\times}}^{2},\quad\mathcal{I}^{\star}(\mathbf{a})=\begin{bmatrix}\mathbf{a}\\ 0\end{bmatrix}$ (1)
We analyze observability in the context of the continuous-time system in Eqs. (2) and (3):
$\dot{\mathbf{x}}(t)$ $=\mathbf{f}(\mathbf{x}(t),\mathbf{u}(t)),\quad\mathbf{x}(0)=\mathbf{x}_{0}$ (2)
$\mathbf{y}(t)$ $=\mathbf{h}(\mathbf{x}(t))+\boldsymbol{\eta}(t),$ (3)
with state space $\mathcal{X}\subset\mathbb{R}^{n\times 1}$, input space $\mathcal{U}\subset\mathbb{R}^{m\times 1}$ and observation space $\mathcal{Y}\subset\mathbb{R}^{p\times 1}$. The deterministic flow $\phi_{t}:\mathcal{X}\rightarrow\mathcal{X}$ describes the solution of the system under a fixed input trajectory $\mathbf{u}(t)$ with $\mathbf{x}(t)=\phi_{t}(\mathbf{x}_{0})$.The measurement noise $\boldsymbol{\eta}(t)$ is modeled as a zero-mean Gaussian process (GP), with a covariance kernel $\mathbf{K}(t,t^{\prime})=E\left[\boldsymbol{\eta}(t)\boldsymbol{\eta}{(t^{\prime})}^{\top}\right]$.
2.2 The Quadrotor Relative Motion Model
The geometry of a pair of cooperatively localizing quadrotor drones is shown in Fig. 1. $\mathbf{p}^{l},\mathbf{p}^{f}\in\mathbb{R}^{3\times 1}$ are the positions of the leader drone and the follower drone, respectively. $\mathbf{r}=\mathbf{R}\left(\mathbf{q}^{f}\right)\left(\mathbf{p}^{l}-\mathbf{p}^{f}\right)\in\mathbb{R}^{3\times 1}$ is the relative position between the leader and the follower. $\mathbf{q}^{l},\mathbf{q}^{f}\in\mathcal{S}^{3}$ are the unit quaternions that represent the orientations of the leader and the follower, respectively. $\mathbf{q}={(\mathbf{q}^{f}})^{-1}\otimes\mathbf{q}^{l}\in\mathcal{S}^{3}$ is the quaternion that represents the relative orientation between the leader and the follower. $\mathbf{v}^{l},\mathbf{v}^{f}\in\mathbb{R}^{3\times 1}$ are the velocities of the leader and the follower, respectively. $\mathbf{v}=\mathbf{R}\left(\mathbf{q}^{f}\right)\left(\mathbf{v}^{l}-\mathbf{v}^{f}\right)\in\mathbb{R}^{3\times 1}$ is the relative velocity between the leader and the follower. $f^{l},f^{f}\in\mathbb{R}^{+}$ are the specific thrusts of the leader and the follower, respectively, and $\boldsymbol{\omega}^{l},\boldsymbol{\omega}^{f}\in\mathbb{R}^{3\times 1}$ are the body rates of the leader and the follower, respectively. The continuous-time relative dynamics of the fleet, which plays the role of $\mathbf{f}$ in Eq. (2), are given as:
$\dot{\mathbf{x}}$ $=\mathbf{f}(\mathbf{x},\mathbf{u})\triangleq\begin{bmatrix}{\mathbf{r}}^{\times}\boldsymbol{\omega}^{l}+\mathbf{v}\\ -\frac{1}{2}\mathcal{I}^{\star}(\boldsymbol{\omega}^{f})\otimes\mathbf{q}+\frac{1}{2}\mathbf{q}\otimes\mathcal{I}^{\star}(\boldsymbol{\omega}^{l})\\ {\mathbf{v}}^{\times}\boldsymbol{\omega}^{f}+f^{l}\mathbf{R}\left(\mathbf{q}\right)\mathbf{1}_{3}-f^{f}\mathbf{1}_{3}\end{bmatrix},\quad\mathbf{u}=\begin{bmatrix}\mathbf{u}^{l}\\ \mathbf{u}^{f}\\ \end{bmatrix}$, (4)
where $\mathbf{x}={\begin{bmatrix}\mathbf{r}^{\top}&\mathbf{q}^{\top}&\mathbf{v}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{10\times 1}$ is the system state, $\mathbf{u}^{l}={\begin{bmatrix}f^{l}&{\boldsymbol{\omega}^{l}}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{4\times 1}$, and $\mathbf{u}^{f}={\begin{bmatrix}f^{f}&{\boldsymbol{\omega}^{f}}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{4\times 1}$ are the inputs corresponding to the leader drone and the follower drone, respectively, and $\mathbf{u}\in\mathbb{R}^{8\times 1}$ is the system input. Equation (4) can be written in a control-affine form $\dot{\mathbf{x}}=\mathbf{f}_{0}(\mathbf{x})+\sum_{i=1}^{m}\mathbf{f}_{i}(\mathbf{x})u_{i}$ as follows:
$\dot{\mathbf{x}}=\underbrace{\begin{bmatrix}\mathbf{v}\\ \mathbf{0}\\ \mathbf{0}\end{bmatrix}}_{\mathbf{f}_{0}}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ \mathbf{R}(\mathbf{q})\mathbf{1}_{3}\end{bmatrix}}_{\mathbf{f}_{1}}f^{l}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{J}^{+}\\ \mathbf{0}\end{bmatrix}}_{\mathbf{f}_{2}}\boldsymbol{\omega}^{l}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ -\mathbf{1}_{3}\end{bmatrix}}_{\mathbf{f}_{3}}f^{f}+\underbrace{\begin{bmatrix}{\mathbf{r}}^{\times}\\ -\mathbf{J}^{-}\\ {\mathbf{v}}^{\times}\end{bmatrix}}_{\mathbf{f}_{4}}\boldsymbol{\omega}^{f},\quad\text{where}\ \mathbf{J}^{\pm}(\mathbf{q})=\begin{bmatrix}q_{4}\mathbf{1}\pm\mathbf{q}_{1:3}^{\times}\\ -\mathbf{q}_{1:3}\end{bmatrix}$. (5)
The discrete-time dynamics based on Eq. (4) can be discretized by integration of RK4 over a timestep $\Delta T$:
$\mathbf{x}_{k+1}=\mathbf{f}_{d}(\mathbf{x}_{k},\mathbf{u}_{k})\triangleq\mathbf{x}_{k}+\texttt{RK4}(\mathbf{f}(\cdot,\mathbf{u}_{k}),\mathbf{x}_{k},\Delta T),$ (6)
2.3 The Inter-drone Distance-only Observation Model
In our CLS, the leader determines its position and velocity through GNSS. The follower is equipped with no global positioning sensors except a UWB sensor, which provides only an inter-drone distance measurement $d=\norm{\mathbf{r}}$. The attitudes of the leader and the follower are measured by their individual IMU and compasses. This information is shared between the leader and follower either by extending the UWB system to transmit richer data alongside range measurements or via Wi-Fi, allowing the CLS to compute the relative pose $\mathbf{q}$. The observation model is as follows:
$\mathbf{y}=\mathbf{h}(\mathbf{x})\triangleq\begin{bmatrix}\frac{1}{2}{d}^{2}\\ \mathbf{q}\end{bmatrix}\in\mathbb{R}^{5\times 1}.$ (7)
where $\frac{1}{2}{d}^{2}$ is used to avoid numerical problems without changing the available information about inter-drone distance.
2.4 The Follower Drone Localization Problem with Range-Only Measurements
For a nonlinear system given by Eqs. (2) and (3), state estimators are typically not optimal, and estimation precision can be improved by trajectory optimization. The classical approach is to optimize some metric $\gamma$, such as the trace, determinant, or minimum eigenvalue, of the EKF posterior covariance matrix $\hat{\mathbf{P}}_{k}$ according to [3, 6] as shown:
$\min_{\mathbf{u}_{1},\ldots,N}\gamma\left(\hat{\mathbf{P}}_{N}\right)$ (8)
$\operatorname{\mathrm{subject\ to\ }}\ \hat{\mathbf{P}}_{k}^{-1}=\mathbf{H}_{k}^{\top}\boldsymbol{\Sigma}\mathbf{H}_{k}+\check{\mathbf{P}}_{k-1},\quad\mathbf{x}_{k+1}=\mathbf{f}_{d}(\mathbf{x}_{k},\mathbf{u}_{k})$ (9)
where $\mathbf{H}_{k}$ is the observation Jacobian and $\check{\mathbf{P}}_{0}$ is the initial prior covariance. However, this approach is not viable for systems where some state variables are not expressed in the observation model. A well-studied example is inertial navigation models augmented by IMU biases [8, 9]. Our CLS is another such example, since the inter-drone velocity $\mathbf{v}$ is not expressed in Eq. (7). In this case, the observation Jacobian contains zero columns at any point of evaluation $\mathbf{x}_{k}$:
$\mathbf{H}_{k}=D\mathbf{h}(\mathbf{x}_{k})=\begin{bmatrix}\mathbf{r}_{k}^{\top}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{1}&\mathbf{0}\end{bmatrix}\in\mathbb{R}^{5\times 10}.$ (10)
[…]
This issue arises because $\mathbf{H}_{k}^{\top}\boldsymbol{\Sigma}\mathbf{H}_{k}$ only accounts for observation in timestep $k$, neglecting the evolution of observation along the flow of the system. Recent approaches to approximating the LOG directly address this through Lie derivatives, which motivates a shift towards the LOG for representing localization quality.
Subsequently, we start by proving that classical local weak observability is attainable by our range-only quadrotor CLS in Section 3 to establish the viability of this system. Then, we address the issue of approximating the LOG efficiently and reasonably through the STLOG in Section 4and use it to develop a control principle in Section 5.
3 Observability Analysis with Range-Only Measurement
In this section, we review the classical nonlinear observability analysis and perform it on our CLS to prove that it satisfies weak local observability. We also introduce a local observability index to prepare for investigating whether weak local observability can be automatically satisfied by observability-aware optimal control. We use the definition from Ref. [18], that a system is observable if and only if the Observability Gramian (OG)
$\mathbf{W}=\int_{0}^{T}\boldsymbol{\Phi}_{t}^{\top}\mathbf{H}^{\top}\mathbf{H}\boldsymbol{\Phi}_{t}dt,$ (11)
is nonsingular. For a nonlinear system, $\mathbf{W}$ has no closed-form expression and varies with the local point of evaluation, so it is referred to as the LOG [7]. In practice, Hermann and Krener’s [19] rank test are used to detect the observability of a dynamical system. For control-affine systems like Eq. (5), the rth-order nonlinear Observability Matrix is:
$\boldsymbol{\mathcal{O}}^{(r)}(\mathbf{x})\triangleq\begin{bmatrix}\vdots\\ DL^{({k})}_{\mathbf{f}_{i},\ldots,j}\mathbf{h}(\mathbf{x})\\ \vdots\end{bmatrix}\quad i,\ldots,j=0,\ldots,m,\quad k=0,\ldots,r$ (12)
[…]
Proposition 3.1.
The range-only quadrotor CLS given by Eqs. (4) and (7) is weakly locally observable, since its 3rd-order observability matrix may be expressed as:
$\boldsymbol{\mathcal{O}}^{(3)}(\mathbf{x})=\left[\begin{array}[]{ccc}{\mathbf{r}}^{\top}&\mathbf{0}_{1\times 4}&\mathbf{0}_{1\times 3}\\ \mathbf{0}_{4\times 3}&\mathbf{1}_{4\times 4}&\mathbf{0}_{4\times 3}\\ \hline\cr{\mathbf{v}}^{\top}&\mathbf{0}_{1\times 4}&{\mathbf{r}}^{\top}\\ \hline\cr\mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 4}&2{\mathbf{v}}^{\top}\\ \hline\cr\mathbf{1}_{3}^{\top}\mathbf{R}{(\mathbf{q})}^{\top}&{\mathbf{r}}^{\top}\mathbf{J}&\mathbf{0}_{1\times 3}\\ \hline\cr-\mathbf{1}_{3}^{\top}&\mathbf{0}_{1\times 4}&\mathbf{0}_{1\times 3}\\ \hline\cr\mathbf{0}_{1\times 3}&\mathbf{0}_{1\times 4}&-2\mathbf{1}_{3}^{\top}\end{array}\right],$ (13)
This matrix attains a rank of 10. Therefore, the system satisfies the observability rank condition.
Proof.
The partitions of $\boldsymbol{\mathcal{O}}^{(3)}(\mathbf{x})$ in Eq. (13) are the nonzero rows of $DL\mathbf{h}$, $DL_{\mathbf{f}_{0}}\mathbf{h}$, $DL_{\mathbf{f}_{0}}^{2}\mathbf{h}$, $DL_{\mathbf{f}_{3}\mathbf{f}_{0}}\mathbf{h}$, $DL_{\mathbf{f}_{1}\mathbf{f}_{0}}\mathbf{h}$, and $DL_{\mathbf{f}_{3}}L_{\mathbf{f}_{0}}^{2}\mathbf{h}$, computed by the MATLAB symbolic toolbox [^5] The relevant code can be found in https://github.com/FSC_Lab/observability_aware_controller/symbolic_algebra. By row reduction, it is found to have a rank of 10.
We can identify that the rank of $\boldsymbol{\mathcal{O}}^{(3)}(\mathbf{x})$ falls below 10 under certain conditions in state $\mathbf{x}$, such as when $\mathbf{r}$ becomes a scalar multiple of $\mathbf{v}$. Although we do not investigate these conditions in detail, we pose the question of whether observability-aware control can direct the state of the system to satisfy weak local observability. To this end, we define local observability index as the minimum order of Lie derivatives necessary for $\boldsymbol{\mathcal{O}}^{(r)}(\mathbf{x},\mathbf{u})$ to attain full rank:
[…]
where:
$\boldsymbol{\mathcal{O}}^{(r)}(\mathbf{x},\mathbf{u})=\begin{bmatrix}\vdots\\ D{L_{\mathbf{f}}^{(k)}\mathbf{h}(\mathbf{x},\mathbf{u})}\\ \vdots\end{bmatrix},\quad k=0,\ldots,r,$ (15)
is another observability matrix, parameterized by control inputs but free of the control-affine requirement, as seen in [8].
Thus, we proved that the quadrotor CLS satisfies weak local observability, and defined the local observability index to help us investigate whether this property is guaranteed under observability-aware control in Section 5.2.
4 The Measure of Observability via STLOG
In this section, we consider the problem of estimating a system’s initial condition from noisy observations along its trajectory, according to the definition of observability, and then derive the STLOG based on our findings.
(Ours) Continuous-time observation process $\mathbf{y}(t)$ with GP noise $\boldsymbol{\eta}(t)$ enables inference without relying on discrete observations (Classical) Discrete observations dictate uncertainty propagation(Ours) IC posteriorinferred at any $t\in[0,T]$$\mathbf{x}_{0}$yyttTT
Figure 2: Visualizing initial state estimation by (our) observation process and (classical) discrete observations
We model an uncertain initial condition as $\mathbf{x}_{0}\sim\mathcal{N}\left(\mathbf{x}_{0}^{*},\check{\mathbf{P}}_{0}\right)$. Across the horizon $[0,T]$, the system receives observations $\mathbf{y}(t)$. Since we are deriving the LOG, a continuous-time quantity, we depart from the classical state estimation approach of analyzing a few discrete observations as illustrated in Fig. 2, and instead model the observation noise $\boldsymbol{\eta}(t)$ as a GP. We represent the estimate for $\mathbf{x}_{0}$ given $\mathbf{y}(t)$ by the posterior distribution $p(\mathbf{x}_{0}\mid\mathbf{y}(t))$, which is factored by Bayes’ rule:
$p(\mathbf{x}_{0}\mid\mathbf{y}(t))=\frac{p(\mathbf{y}(t)\mid\mathbf{x}_{0})}{p(\mathbf{y}(t))}p(\mathbf{x}_{0}),$ (16)
The key challenge in this problem is that the likelihood ratio ${p(\mathbf{y}(t)|\mathbf{x}_{0})}/{p(\mathbf{y}(t))}$ is between infinite-dimensional distributions, which is less tractable for analysis and computation than the elementary, finite-dimensional case. We outline these difficulties heuristically and defer a full rigorous treatment using Gaussian measure theory to Appendix A.
We consider the covariance kernel $\mathbf{K}(t,t^{\prime})=E\left[\boldsymbol{\eta}(t)\boldsymbol{\eta}{(t^{\prime})}^{\top}\right]$ of GP $\boldsymbol{\eta}(t)$, and we consider the map from an initial condition to a function in time of motion-coupled observations to be the system response $\mathcal{A}:\mathcal{X}\rightarrow C^{\infty}([0,T],\mathbb{R}^{p})$, defined as:
$\left(\mathcal{A}(\mathbf{x}_{0})\right)(t)=\mathcal{A}_{t}(\mathbf{x}_{0})=\mathbf{h}\left(\phi_{t}(\mathbf{x}_{0})\right)=\mathbf{h}\left(\mathbf{x}(t)\right),$ (17)
Subject to the system model in Eq. (2) and (3), the conditional observation $\mathbf{y}(t)|\mathbf{x}_{0}$ is also a GP, with mean $\mathcal{A}_{t}(\mathbf{x}_{0})$ and covariance $\mathbf{K}(t,t^{\prime})$. This enables us to consider a subproblem of computing the likelihood ratio between two initial conditions, ${p(\mathbf{y}(t)|\mathbf{x}_{0})}/{p(\mathbf{y}(t)|\mathbf{x}_{1})}$. Recall from Mercer’s theorem in Ref. [20] that the covariance kernel $\mathbf{K}(t,t^{\prime})$, under suitable assumptions, defines an integral operator $\mathcal{K}$ in the Hilbert space $L^{2}([0,T],\mathbb{R}^{p})$ of vector-valued square-integrable functions on $[0,T]$. The operator $\mathcal{K}$ has a complete set of orthonormal eigenfunctions $\left\{\mathbf{v}_{n}(t)\right\}_{n=0}^{\infty}$ with respect to the inner product $\langle\cdot,\cdot\rangle$ in $L^{2}([0,T],\mathbb{R}^{p})$}, with corresponding eigenvalues $\left\{\lambda_{n}\right\}_{n=0}^{\infty}$ such that $(\mathcal{K}\mathbf{v}_{n})(t)=\lambda_{n}\mathbf{v}_{n}(t)$.
Let $y_{n}\lvert_{i}=\langle\mathbf{v}_{n}(t),\mathbf{y}(t)\rangle|\mathbf{x}_{i}$ be the component of $\mathbf{y}(t)|\mathbf{x}_{i}$ in the eigenfunction basis defined by $\mathcal{K}$. Each $y_{n}\lvert_{i}$ is then a one-dimensional Gaussian variable, with mean $\langle\mathbf{v}_{n}(t),\mathcal{A}_{t}(\mathbf{x}_{i})\rangle$ and variance $\lambda_{n}$. Moreover, $y_{n}\lvert_{i}$ and $y_{m}\lvert_{i}$ are independent of $n\neq m$. An elementary attempt to evaluate ${p(\mathbf{y}(t)|\mathbf{x}_{0})}/{p(\mathbf{y}(t)|\mathbf{x}_{1})}$ is to treat it as the infinite product of the component-wise likelihood ratios:
$\frac{p(\mathbf{y}(t)|\mathbf{x}_{0})}{p(\mathbf{y}(t)|\mathbf{x}_{1})}\overset{heuristic}{=}\frac{\Pi_{n=0}^{\infty}p(y_{n}\lvert_{0})}{\Pi_{n=0}^{\infty}p(y_{n}\lvert_{1})}\overset{heuristic}{=}\Pi_{n=0}^{\infty}\frac{p(y_{n}\lvert_{0})}{p(y_{n}\lvert_{1})},$ (18)
where we have annotated the heuristic nature of the relation between each expression. The heuristic expressions in Eq. (18) elicit questions regarding:
  1. 1.
    Whether the ‘joint probability distribution’ $\Pi_{n=0}^{\infty}p(y_{n}\lvert_{i})$ exists meaningfully; and if it does, whether its samples belong to a sufficiently regular function space, such as $L^{2}([0,T],\mathbb{R}^{p})$}.
  2. 2.
    Whether the infinite product of likelihood ratios $\Pi_{n=0}^{\infty}({p(y_{n}\lvert_{0})}/{p(y_{n}\lvert_{1})})$ converges.
We address these questions rigorously in Appendix A, but state the main result for proceeding with the initial condition estimation problem in Proposition 4.1.
Proposition 4.1.
Under suitable assumptions on the system response $\mathcal{A}$ in Eq. (17) and the covariance operator $\mathcal{K}$, the likelihood ratio ${p(\mathbf{y}(t)|\mathbf{x}_{0})}/{p(\mathbf{y}(t))}$ can be expressed in closed-form as:
$\frac{p(\mathbf{y}(t)|\mathbf{x}_{0})}{p(\mathbf{y}(t))}=C(\mathbf{y}(t))\exp\left(\langle\mathbf{y}(t),\mathcal{K}^{-1}\mathcal{A}_{t}(\mathbf{x}_{0})\rangle-\frac{1}{2}\langle\mathcal{A}_{t}(\mathbf{x}_{0}),\mathcal{K}^{-1}\mathcal{A}_{t}(\mathbf{x}_{0})\rangle\right).$ (19)
where $\langle\cdot,\cdot\rangle$ denotes the inner product on $L^{2}([0,T],\mathbb{R}^{p})$}.
Proof.
The conditions for Eq. (19) to be derived is given in Proposition A.2 and Corollary A.2.1
Next, we introduce a practical noise model, which, when paired with the assumption that the system response $\mathcal{A}:\mathcal{X}\rightarrow C^{\infty}([0,T],\mathbb{R}^{p})$ in Eq. (17) is continuous with respect to the Fréchet topology on $C^{\infty}([0,T],\mathbb{R}^{p})$666 A characterization of the Fréchet topology on $C^{\infty}([0,T],\mathbb{R}^{p})$ is as the metric topology of $d(\mathbf{f}(t),\mathbf{g}(t))=\sum_{k=0}^{\infty}2^{-k}\frac{\rho_{k}(\mathbf{f}-\mathbf{g})}{1+\rho_{k}(\mathbf{f}-\mathbf{g})}$ [21] where $\rho_{k}(\mathbf{f}-\mathbf{g})={\sup}_{t\in[0,T]}\norm{\frac{d^{k}}{dt^{k}}(\mathbf{f}(t)-\mathbf{g}(t))}.$ In particular, all linear systems will have a continuous system response, owing to the existence of an explicit form $\mathcal{A}_{t}(\mathbf{x}_{1})-\mathcal{A}_{t}(\mathbf{x}_{0})=\mathbf{M}(t)\left(\mathbf{x}_{1}-\mathbf{x}_{0}\right)$, for some analytically known smooth matrix-valued function $\mathbf{M}(t)$ on $[0,T]$ [18]. , satisfies the prerequsites of Proposition 4.1. We achieve this by specifying the covariance operator $\mathcal{K}$ spectrally, using the orthonormal Legendre polynomials ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ on $[0,T]$. The detailed rationale is explained in Corollary A.2.1 and Example A.1.
Definition 4.1 (Band-limited white noise model with pink tail).
$\boldsymbol{\eta}(t)=\frac{1}{\sqrt{N_{c}}}\sum_{n=0}^{N_{c}-1}\boldsymbol{\chi}_{n}e_{n}(t)+\sum_{n=N_{c}}^{\infty}\sqrt{a_{n}n^{-\alpha}}\boldsymbol{\chi}_{n}e_{n}(t),$ (20)
where
  1. 1.
    $N_{c}>1$: Bandwidth parameter, signifying the number of ‘active’ Legendre components
  2. 2.
    $\alpha>1$: High-frequency decay exponent
  3. 3.
    ${\{\boldsymbol{\chi}_{n}\}}_{0}^{\infty}$: A sequence of independent, vector-valued spatial noise on each Legendre component, each following the Gaussian distribution $\mathcal{N}(0,\boldsymbol{\Sigma})$, where $\boldsymbol{\Sigma}$ is a positive-definite covariance matrix.
  4. 4.
    ${\{a_{n}\}}_{N_{c}}^{\infty}$: A sequence of tail coefficients. Positive constants $a_{-},a_{+}$ must exist such that $a_{n}\in[a_{-},a_{+}]$ for all $n$.
Definition 4.1 approximately defines a GP with $N_{c}$ ‘active’ components, each following the prescribed Gaussian distribution $\mathcal{N}(0,\boldsymbol{\Sigma})$ and with equal energy $1/N_{c}$, plus a power-law decay at frequencies above $1/\delta t=N_{c}/T$. More precisely, let ${\left\{\sigma_{i}\right\}}_{i=1}^{p}$ and ${\left\{\mathbf{v}_{i}\right\}}_{i=1}^{p}$ be the sets of eigenvalues and orthonormal eigenvectors for $\boldsymbol{\Sigma}$, respectively. Then $\left\{\mathbf{v}_{i}e_{n}(t)\right\}$} is a complete set of orthonormal eigenfunctions for the covariance operator $\mathcal{K}$, with eigenvalues:
$\mathcal{K}\left(\mathbf{v}_{i}e_{n}(t)\right)=\begin{cases}\frac{1}{N_{c}}\sigma_{i}\mathbf{v}_{i}e_{n}(t)&n<N_{c},\\ a_{n}n^{-\alpha}\sigma_{i}\mathbf{v}_{i}e_{n}(t)&n\geq N_{c}.\end{cases}$ (21)
We follow Example A.1 to check that $\mathcal{K}$ in Eq. (21) satisfies the relevant hypotheses in Corollary A.2.1 and apply Proposition 4.1 to express the likelihood ratio as Eq. (19). Then, we further simplify it in Proposition 4.2.
Proposition 4.2.
Using the noise model in Definition 4.1, for sufficiently large $N_{c}$, the likelihood ratio in Eq. (19) at a given $\mathbf{x}_{0}$ and a given $\mathbf{y}(t)$ can be written as
$\frac{p(\mathbf{y}(t)|\mathbf{x}_{0})}{p(\mathbf{y}(t))}=C(\mathbf{y}(t))\exp\left(\frac{N_{c}}{T}\left(\int_{0}^{T}\mathbf{y}^{\top}(t)\boldsymbol{\Sigma}^{-1}\mathcal{A}_{t}(\mathbf{x}_{0})-\frac{1}{2}\mathcal{A}_{t}^{T}(\mathbf{x}_{0})\boldsymbol{\Sigma}^{-1}\mathcal{A}_{t}(\mathbf{x}_{0})\ dt\right)+{O}(N_{c}^{-k})\right),$ (22)
for all $k>0$.
Proof.
Since $\mathcal{A}(\mathbf{x}_{0})(t)$ is smooth in $t$, its $n$-th Fourer–Legendre coefficient is ${O}(n^{-k})$ for all $k>0$ (see Example A.1). Using the explicit expression for $\mathcal{K}$ in Eq. (21), we compute that, for sufficiently large $N_{c}$,
$\left(\mathcal{K}^{-1}\mathcal{A}(\mathbf{x}_{0})\right)(t)=N_{c}\boldsymbol{\Sigma}^{-1}\mathcal{A}_{t}(\mathbf{x}_{0})+{O}(N_{c}^{-k}),$ (23)
for all $k>0$. Substituting Eq. (23) into Eq. (19) gives the desired result. ∎
In the following analysis, we shall neglect the ${O}(N_{c}^{-k})$ terms in Eq. (22). We now describe the structure of the posterior logarithmic likelihood $\log p(\mathbf{x}_{0}\mid\mathbf{y}(t))$ as a function of $\mathbf{x}_{0}$, up to constants $C_{1},C_{2}$ independent of $\mathbf{x}_{0}$. We express this equivalent class using Eqs. (16)and (22):
$\log p(\mathbf{x}_{0}|\mathbf{y}(t))+C=\log\frac{p(\mathbf{y}(t)|\mathbf{x}_{0})}{p(\mathbf{y}(t))}+\log p(\mathbf{x}_{0})+C=\log\left(\frac{p(\mathbf{y}(t)|\mathbf{x}_{0})}{p(\mathbf{y}(t))}{\left(\frac{p(\mathbf{y}(t)|\mathbf{x}^{*}_{0})}{p(\mathbf{y}(t))}\right)}^{-1}\right)+\log p(\mathbf{x}_{0})+C^{\prime}$
$=\frac{N_{c}}{T}\int_{0}^{T}\left({(\mathbf{y}(t)-\mathcal{A}_{t}(\mathbf{x}_{0}^{*}))}^{T}\boldsymbol{\Sigma}^{-1}(\mathcal{A}_{t}(\mathbf{x}_{0})-\mathcal{A}_{t}(\mathbf{x}_{0}^{*}))-{(\mathcal{A}_{t}(\mathbf{x}_{0})-\mathcal{A}_{t}(\mathbf{x}_{0}^{*}))}^{T}\frac{1}{2}\boldsymbol{\Sigma}^{-1}(\mathcal{A}_{t}(\mathbf{x})-\mathcal{A}_{t}(\mathbf{x}_{0}^{*}))\right)dt$
$\quad-\frac{1}{2}{(\mathbf{x}-\mathbf{x}_{0}^{*})}^{T}\check{\mathbf{P}}_{0}^{-1}(\mathbf{x}-\mathbf{x}_{0}^{*}),$ (24)
where we introduced the mean of the initial condition $\mathbf{x}_{0}^{*}$ as a reference point and expressed the likelihood ratio relative to it to eliminate the intractable marginal likelihood $p(\mathbf{y}(t))$. Since $\mathbf{x}_{0}^{*}$ is fixed, its associated likelihood ratio $\frac{p(\mathbf{y}(t)|\mathbf{x}^{*}_{0})}{p(\mathbf{y}(t))}$ contributes only an additive constant, which is absorbed into the change from $C$ to $C^{\prime}$. This preserves the equivalence class of the posterior up to terms independent of $\mathbf{x}_{0}$, while allowing the remaining terms to be expressed in a tractable quadratic form.
To simplify Eq. (4), we linearize around the prior expectation $\mathbf{x}_{0}^{*}$. Such a linearization is valid in the following scaling regime: for some scaling parameter ${\epsilon}\ll 1$,
$\delta\mathbf{x}=\mathbf{x_{0}}^{*}-\mathbf{x}_{0}\thicksim{\epsilon},\quad\boldsymbol{\Sigma}\thicksim{\epsilon}^{2},\quad\check{\mathbf{P}}_{0}\thicksim{\epsilon}^{2},$ (25)
[…]
$\log p(\delta\mathbf{x}|\mathbf{y}(t))+C=\frac{N_{c}}{T}\mathbf{b}^{\top}\delta\mathbf{x}+\frac{1}{2}\delta\mathbf{x}^{\top}\left(\check{\mathbf{P}}_{0}^{-1}+\frac{N_{c}}{T}\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}\right)\delta\mathbf{x},$ (26)
to leading order of $\epsilon$}, where
$\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}=\int_{0}^{T}D\mathcal{A}_{t}^{\top}\boldsymbol{\Sigma}^{-1}D\mathcal{A}_{t}dt,\quad\mathbf{b}=\int_{0}^{T}D\mathcal{A}_{t}^{\top}\boldsymbol{\Sigma}^{-1}\mathbf{z}(t)dt,$ (27)
and $\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}$ corresponds to the LOG with respect to the metric $\boldsymbol{\Sigma}^{-1}$ in the observation space $\mathcal{Y}$.
Remark.
In the special case $\boldsymbol{\Sigma}=\mathbf{1}$ and $\mathcal{A}_{t}=\mathbf{h},\left(\phi_{t}(\mathbf{x}_{0})\right)$ is smooth, substituting the classical system response map $D\mathcal{A}_{t}=\mathbf{H}\boldsymbol{\Phi}_{t}$ into Eq. (27) gives the traditional definition of continuous time OG in Eq. (11).
Under this linear approximation, $\delta\mathbf{x}$ becomes a normally distributed random variable, with expectation and covariance
${E}[\delta\mathbf{x}|\mathbf{y}(t)]=\mathbb{V}[\delta\mathbf{x}|\mathbf{y}(t)]\left(\frac{N_{c}}{T}\mathbf{b}\right),\quad\mathbb{V}[\delta\mathbf{x}|\mathbf{y}(t)]={\left(\frac{N_{c}}{T}{\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}}+\check{\mathbf{P}}_{0}^{-1}\right)}^{-1}.$ (28)
We conclude that in a small-noise regime where the scaling in Eq. (25) holds, the posterior of the initial condition estimate under a given observation path $\mathbf{y}(t)$ at times $[0,T]$ is approximately Gaussian $\mathbf{x}_{0}\sim\mathcal{N}(\hat{\mathbf{x}}_{0},\hat{\mathbf{P}}_{0})$, where the posterior mean and covariance are given by
$\hat{\mathbf{x}}_{0}-\mathbf{x}_{0}^{*}=\frac{1}{\delta t}\hat{\mathbf{P}}_{0}\mathbf{b},\quad\hat{\mathbf{P}}_{0}^{-1}=\frac{1}{\delta t}{\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}}+\check{\mathbf{P}}_{0}^{-1},$ (29)
where $\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}$ and $\mathbf{b}$ given in Eq. (27), and $\delta t=T/N_{c}$ is the correlation timescale of the noise term $\boldsymbol{\eta}(t)$. Moreover, the relationship between prior and posterior covariances $\check{\mathbf{P}}_{0},\hat{\mathbf{P}}_{0}$ in Eq. (29) is still valid if the prior distribution of $\mathbf{x}_{0}$ has a different prior mean $\check{\mathbf{x}}_{0}$, provided that the scaling $\check{\mathbf{x}}_{0}-\mathbf{x}^{*}_{0}\sim\epsilon$ holds, because the linearization argument for the log-posterior in Eq. 4 remains valid. In this regime, the minimum precision (inverse covariance) of the posterior distribution satisfies the following lower bound:
${\lambda_{\max}(\hat{\mathbf{P}}_{0})}^{-1}\geq\frac{\lambda_{\min}(\mathbf{W}_{\boldsymbol{\Sigma}^{-1}})}{\delta t}+{\lambda_{\max}(\check{\mathbf{P}}_{0})}^{-1},$ (30)
where $\lambda_{\min/\max}$ denotes the minimum/maximum eigenvalue of the relevant matrix.
Equation (30) enables us to evaluate arbitrarily long prospective trajectories based on their contribution to improving initial condition estimation precision. To use this ability in a control context, we derive the STLOG as an approximation of the LOG based on $\mathcal{A}_{t}$ from Eq. (17). This STLOG approximation is based on the following conditions
  1. 1.
    The observation horizon $T$ is sufficiently short, but not shorter than the correlation timescale $\delta t$ of the noise $\boldsymbol{\eta}(t)$ in Definition 4.1, so that $\delta t\ll T\ll 1$;777 The STLOG approximation to the LOG is independent of the correlation timescale $\delta t$ per se, but the ordering $\delta t\ll T$ is required for the LOG to emerge from Bayesian state estimation (c.f. Proposition 4.2).
  2. 2.
    $\mathbf{u}(t)$ is constant for $t\in[{0},{T})$, so that our system dynamics in Eq. (2) is autonomous within this horizon.
The STLOG is fundamentally based on the expansion of the system response, defined as $\mathcal{A}_{t}(\mathbf{x}_{0})=\mathbf{h}(\mathbf{x}(t))$ in Eq. (17) as a Taylor series in $t$ about $t=0$. Differentiating Eq. (17) with respect to $t$ under the conditions that $\mathbf{u}(t)$ is constant and $\mathbf{f}$ is time-independent gives the derivatives:
$\frac{d}{dt}\mathcal{A}_{t}(\mathbf{x}_{0})=\left(L_{\mathbf{f}}\mathbf{h}\right)(\mathbf{x}(t)),\quad\frac{d^{j}}{dt^{j}}\mathcal{A}_{t}(\mathbf{x}_{0})=\left(L_{\mathbf{f}}^{j}\mathbf{h}\right)(\mathbf{x}(t)),$ (31)
where $L_{\mathbf{f}}=\mathbf{f}\cdot\nabla$ is the Lie derivative on functions [^8] This is consistent with the definition of Lie derivatives in differential geometry (see, e.g. [22]) if we consider $\mathbf{h}$ as a $p$-tuple of scalar functions in the manifold $\mathcal{X}$, whence no contractions involving $D\mathbf{f}$ are necessary. Note, however, that $D\mathbf{h}$ is considered as a $p$-tuple of one-forms on $\mathcal{X}$, and correspondingly $L_{\mathbf{f}}D\mathbf{h}=\mathbf{f}\cdot\nabla(D\mathbf{h})+(D\mathbf{h})(D\mathbf{f})$. The usual identity $L_{\mathbf{f}}D\mathbf{h}=DL_{\mathbf{f}}\mathbf{h}$ holds under this interpretation.. We express $\mathcal{A}_{t}(\mathbf{x}_{0})$ in terms of the derivatives in Eq. (31) taken at $t=0$ to get:
$\mathcal{A}_{t}(\mathbf{x}_{0})=\sum_{j=0}^{\infty}\frac{t^{j}}{j!}\left(L_{\mathbf{f}}^{j}\mathbf{h}\right)(\mathbf{x}_{0}),$ (32)
which can be differentiated with respect to the state to give:
$D\mathcal{A}_{t}=\sum_{j=0}^{\infty}\frac{t^{j}}{j!}D\left(L_{\mathbf{f}}^{j}\mathbf{h}\right).$ (33)
Substituting the series in Eq. (33) into $\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}$ in (27) and integrating yields the STLOG in Definition 4.2.
Definition 4.2.
The order-$r$ short-term local observability Gramian (STLOG) is the matrix
$\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}=\sum_{i,j=0}^{r}\frac{{T}^{i+j+1}}{(i+j+1)i!j!}D{\left(L_{\mathbf{f}}^{i}\mathbf{h}\right)}^{\top}\boldsymbol{\Sigma}^{-1}D\left(L_{\mathbf{f}}^{j}\mathbf{h}\right).$ (34)
which would be evaluated at $(\mathbf{x}^{*},\mathbf{u};T)$, respectively, the initial condition, control input, and observation time. As such, the STLOG is a novel point-wise representation of observability, in contrast to the traditional LOG, which requires integration over an entire trajectory. It enables local assessment of how the evolution of the system under the current control input contributes to the observability of the current operating state, and aligns with our goal of leveraging observability within predictive control frameworks designed in Section 5.
5 Observability-aware control
We begin by deriving our observability-aware controller, the OPC, in Section 5.1. We then investigate whether the resulting control strategy satisfies local weak observability in Section 5.2 to address the question raised in Section 3.
5.1 OPC
The goal of observability-aware control is to design a control strategy that refines the estimates of $N$ states $\mathbf{x}(t)$ in any $T_{1},\ldots,T_{k},\ldots,T_{N}$, separated by $\Delta T$ by a discretization scheme. We first propose a law that relates the prior and posterior covariance of state estimation at every $T_{k}$ in a trajectory, by extending Eq. (29) in Proposition 5.1.
Assumption 1.
We assume the inputs $\mathbf{u}(t)$ are piecewise-constant within each stage ${T_{k}},{T_{k}+\Delta T})$, and moreover that the timescales follow the scaling $\delta t\ll\Delta T\ll T$, so we can employ STLOG in Eq. (34) to approximate the LOG.
Assumption 2.
We assume that, for some scale parameter $\epsilon\ll 1$, both the state estimation covariances and $\boldsymbol{\Sigma}$ scale as $\epsilon^{2}$, as in Eq. (25), and that displacements $\delta\mathbf{x}_{t}=\mathbf{x}(t)-\mathbf{x}^{*}(t)$ from the deterministic state $\mathbf{x}^{*}(t)$ scale as $\epsilon$, so that the linearized result Eq. (29) remains valid.
Proposition 5.1.
Under Assumptions 1 and 2, the prior and posterior covariances at step $k$ are related by:
$\hat{\mathbf{P}}_{T_{k}}^{-1}=\frac{\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}\left(\mathbf{x}^{*}(T_{k}),\mathbf{u}(T_{k})\right)}{\delta t}+\check{\mathbf{P}}_{T_{k}}^{-1},$ (35)
where $\check{\mathbf{P}}_{T_{k}}$ is the covariance at time $T_{k}$ given only past observations, and $\hat{\mathbf{P}}_{T_{k}}$ is the covariance at time $T_{k}$ given additional observations between $T_{k}$ and $T_{k+1}$, and $\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}$ is the rth-order STLOG, with observation time $\Delta T$. In particular, an analogous bound to Eq. (30) holds: The posterior precision due to the observations in ${T_{k}},{T_{k+1}})$ is bounded below by
${\lambda_{\max}^{-1}\left(\hat{\mathbf{P}}_{T_{k}}\right)}\geq\frac{\lambda_{\min}\left(\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}\left(\mathbf{x}^{*}(T_{k}),\mathbf{u}(T_{k})\right)\right)}{\delta t}+{\lambda_{\max}^{-1}\left(\check{\mathbf{P}}_{T_{k}}\right)}.$ (36)
Proof.
By analogy to the argument used to derive Eq. (29), we perform linearization about $\mathbf{x}^{*}(t)=\phi_{t}(\mathbf{x}^{*}_{0})$. By treating each prior mean $\check{\mathbf{x}}(T_{k})$ as an initial condition with stage-wise observation timestep $\Delta T$ and linearizing about $\mathbf{x}^{*}(T_{k})$ in the small noise regime in Eq. (25), we apply Eq. 29 to get the following evolution laws for the prior and posterior covariances at each $T_{k}$:
$\check{\mathbf{P}}_{T_{k+1}}=\boldsymbol{\Phi}_{T_{k},T_{k+1}}\hat{\mathbf{P}}_{T_{k}}\boldsymbol{\Phi}_{T_{k},T_{k+1}}^{\top},\quad\hat{\mathbf{P}}_{T_{k}}^{-1}=\frac{\mathbf{W}_{\boldsymbol{\Sigma}^{-1}}\left[\mathbf{x}^{*}(T_{k}),\mathbf{u}:[{T_{k}},{T_{k+1}})\rightarrow\mathcal{U}\right]}{\delta t}+\check{\mathbf{P}}_{T_{k}}^{-1}.$ (37)
where $\boldsymbol{\Phi}_{T_{k},T_{k+1}}=D\left(\boldsymbol{\phi}_{T_{k+1}}\boldsymbol{\phi}^{-1}_{T_{k}}\right)$ is the state transition map from time $T_{k}$ to $T_{k+1}$.999 Similar evolution laws relating the estimation means can also be obtained, but we omit their derivation here since they do not influence the evolution of covariances once the linear approximation around $\mathbf{x}^{*}(t)=\phi_{t}(\mathbf{x}^{*}_{0})$ is admitted, as in Kalman filtering theory for linear systems. Substituting the STLOG approximation in Eq. (34) into Eq. (37) leads to Eq. (35).
Our observability-aware control principle is to collectively maximize the lower bounds to precision improvements, embodied by the minimum eigenvalue of the STLOG in Eq. (36), along a trajectory. This is represented by the objective:
$V(\mathbf{x}^{*}(T_{0,\ldots,N-1}),\mathbf{u}(T_{0,\ldots,N-1}))=\sum_{k=0}^{N-1}{\lambda_{\min}\left(\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}\left(\mathbf{x}^{*}(T_{k}),\mathbf{u}(T_{k})\right)\right)},$ (38)
The observability-predictive controller (OPC) optimizes $V$ under a receding horizon framework with the finite horizon taken to be $N=20$ in our software implementation:
$\min_{\mathbf{u}_{0,\ldots,N-1}}$ $\frac{1}{V(\mathbf{x}_{k},\mathbf{u}_{k})+c}$ (39)
$\operatorname{\mathrm{subject\ to\ }}\$$\begin{cases}\mathbf{x}_{k+1}=\mathbf{f}_{\mathrm{d}}(\mathbf{x}_{k},\mathbf{u}_{k}),\ \mathbf{x}_{0}=\mathbf{x}_{cl},\\ u_{\min,i}\leq\mathbf{u}_{k,i}\leq u_{\max,i},\\ d_{\min}\leq d_{k}\leq d_{\max}.\end{cases}$ (40)
where we adopt the shorthands $\mathbf{x}(T_{k})=\mathbf{x}_{k}$, $\mathbf{u}(T_{k})=\mathbf{u}_{k}$. $\mathbf{x}_{cl}$ denotes the current state estimate provided by the CLS, $c$ is a small regularization constant which we take to be $c=10^{-6}$, $d_{k}$ is the inter-drone distance as part of $\mathbf{x}_{k}$, and $d_{\min},d_{\max}$ are the minimum and maximum bounds on $d_{k}$, respectively.
5.2 Satisfaction of local weak observability under OPC control
Next, we address whether the optimal trajectories produced by the OPC inherently satisfy local weak observability. We recover the discrete rank-testing notion of observability by proposing that the asymptotics of the summand $\lambda_{\min}\left(\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}\left(\mathbf{x},\mathbf{u};\Delta T\right)\right)$ in Eq. (38) as $\Delta T\to 0$ detects the local observability index $r_{*}(\mathbf{x},\mathbf{u})$ in Eq. (14).
Proposition 5.2.
Considering a fixed operating point $(\mathbf{x},\mathbf{u})$,
  1. 1.
    If $r<r_{*}(\mathbf{x},\mathbf{u})$, then $\lambda_{\min}(\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}(\mathbf{x},\mathbf{u};\Delta T))=0$,
  2. 2.
    If $r\geq r_{*}(\mathbf{x},\mathbf{u})$, then for all sufficiently small $\Delta T$,
    $\lambda_{\min}(\mathbf{W}^{(r)}_{\boldsymbol{\Sigma}^{-1}}(\mathbf{x},\mathbf{u};\Delta T))=C(\mathbf{x},\mathbf{u};\Delta T)\Delta T^{2r_{*}+1},$ (41)
    where $C(\mathbf{x},\mathbf{u};\Delta T)$ is bounded between two positive quantities $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ independent from $\Delta T$.
Proof.
This proposition is proved in detail in Appendix B. ∎
Proposition 5.2 suggests that in the OPC, stages yielding low $r_{*}(\mathbf{x},\mathbf{u})$ make asymptotically larger contributions to the objective compared to stages yielding high $r_{*}(\mathbf{x},\mathbf{u})$. Thus, optimized trajectories drive the state $\mathbf{x}$ to avoid regions with high $r_{*}$ and favor regions where $r_{*}$ attains the minimum value. We derive a lower bound for this minimum $r_{*}$ in Proposition C.1, and apply it to our quadrotor CLS in Corollary C.1.1 to obtain $r_{*}\geq 5$, providing us a principled guideline for selecting $r$ for the STLOG in the OPC.
This result also connects our OPC to previous work optimizing metrics of the observability matrix [15]. $C_{1},C_{2}$ in Proposition 5.2 are shown by Propositions B.4 and B.7 to be:
$C_{1}=\alpha\ \sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right),\quad C_{2}=\beta\ \sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\ker{\boldsymbol{\mathcal{O}}^{(r_{*}-1)}}}\right),$ (42)
where $\sigma_{\min}$ denotes the minimum singular value, $|$ denotes the restriction to a subspace, and $\alpha,\beta$ are numerical constants. We identify that a new metric of observability $C^{\prime}$, formed by taking a weighted average of $C_{1}$ and $C_{2}$, can be heuristically substituted into $C$ in Eq. (41) and the objective formed by summing this quantity over a trajectory:
$\mathcal{V}(\mathbf{x}^{*}(T_{0,\ldots,N-1}),\mathbf{u}(T_{0,\ldots,N-1}))\simeq\sum_{k=0}^{N-1}C^{\prime}\left(\mathbf{x}^{*}(T_{k}),\mathbf{u}(T_{k})\right)\Delta T^{2r_{*}\left(\mathbf{x}^{*}(T_{k}),\mathbf{u}(T_{k})\right)+1},$ (43)
has the same physical meaning as the objective in Eq. (38), in that for the same value of $r_{*}$, stages with a higher metric of observability $C^{\prime}$ will contribute more to $\mathcal{V}$, leading to optimal observability trajectories under the metric $C^{\prime}$.
Overall, we met our design goals, as the OPC maximizes the worst-case precision improvement along a trajectory. We also answered the question raised in Section 3, showing that the OPC is compatible with the implementation of weak local observability, then used the result to tailor the selection of $r$ in the OPC to our range-only quadrotor CLS.
6 Experimental Validation
[…]
Figure 3: The overall flow chart of the system.
We evaluate our observability-aware controller on CL-aided quadrotors in two stages. In Section 6.1, we describe our Monte Carlo simulations, modelled after our previous work [6], of an EKF-based CL system configured according to Fig. 3. We evaluate the localization performance of this system when guided by our controller compared to two baseline trajectories. In Section 6.2, we describe our quadrotor flight experiments using the same EKF-based CLS. Collectively, the simulations and experiments assess the performance of observability-aware control under realistic sensing and quadrotor dynamics, with a focus on estimation accuracy and precision.
To compute the STLOG in Eq. (39) in software, we use jax [23] to carry out automatic differentiation and evaluate the Lie derivative of our system defined by Eq. (4) up to r=5r=5101010Python implementation of the OPC can be found inhttps://github.com/FSC_Lab/observability_aware_controller. However, since there is limited upstream support for jax on board computers with the ARM architecture, we cannot compute STLOG in real-time on board. Instead, we use the OPC to compute control inputs online in simulations, but run it offline as a sliding-window trajectory planner to pre-plan trajectories for the quadrotors in experiments.
6.1 Simulation
In simulations, the leader follows a straight trajectory, while the follower performs three test cases:
Case 1
A uniform and straight trajectory, planned ahead-of-time by a minimum-snap trajectory planner,
Case 2
An artificially perturbed, zigzagging trajectory, likewise planned ahead-of-time, and
Case 3
An observability-aware trajectory performed with the OPC computing control inputs in real-time.
Figure 4: Trajectories executed by the quadrotor team in simulations; Squares and circles mark starting and end positions, respectively
We use a range-only CLS based on an EKF as configured in Table 2 to estimate the states of both quadrotors, so that estimation performance can be compared across the three test cases. We implemented realistic noise models for both control inputs and observations received by the CLS. The simulations are run for 120 seconds, with 50 Monte Carlo trials per test case. The same noise regime is applied to all trials.
Table 2: Configuration of CLS in simulation
Parameter Value Notes
Process noise $\sigma^{2}\left(f^{\{f,l\}}\right)$ 0.05(unitless)0.05\ ($\mathrm${unitless}) Thrust model uncertainty
$\sigma^{2}\left(\boldsymbol{\omega}^{\{f,l\}}\right)$ $1.49\cdot 10^{-5}\ \mathrm{rad}^{2}/s^{2}$
[…]
$\sigma^{2}\left(\boldsymbol{\phi}\right)$ $3.594\cdot 10^{-6}\ \mathrm{rad}^{2}$ Per-rotation axis
Figure 4 shows the trajectories trialed in simulation. Our optimized trajectories in Case 3 exhibit visually distinct behaviors — including periodic orbiting patterns — to deliberately excite weakly observable directions. These motions emerge from the OPC without specific hints or deliberate configuration, reproducing the helical [24] and orbiting [17] behavior of observability-optimal trajectories in previous works.
(a) X-axis position error; Horizontal lines show RMS error
(b) Estimated X-axis uncertainty ($3\sigma$ envelope)
(c) Y-axis position error; Horizontal lines show RMS error
(d) Estimated Y-axis uncertainty ($3\sigma$ envelope)
(e) Z-axis position error; Horizontal lines show RMS error
(f) Estimated Z-axis uncertainty ($3\sigma$ envelope)
Figure 5: Results of Monte Carlo simulations of EKF-based CLS along trajectories generated in test cases 1, 2, and 3 (ours). Position estimation error, RMSE error, and uncertainty ($3\sigma$ envelope) are shown along each axis
Figures 5a to 5f visualize the estimation performance of the EKF-based CLS in simulations. The straight trajectory yields in Case 1 a steadily diverging error and high covariance due to a lack of perception- or observability-aware excitation, in line with long-standing conclusions about the pathologies of uniform straight trajectories [3]. The zigzag trajectory in Case 2 is dynamically rich but estimation-unware. It enabled the CLS to occasionally reduce position estimation error at the cost of strong spikes in error that drove up the RMS error. This confirms that simply exciting position space, adding curvature, or desynchronizing follower motion relative to the leader does not guarantee informative sensor observations. Geometrically rich but non-optimized minimum-snap trajectories may also worsen estimation due to poor alignment with observable modes and increased tracking difficulty. In contrast, the observability-optimal (OO) trajectory explicitly excites low-observability directions and maintains low positioning error and tight estimator confidence bounds.
Table 3: Positioning accuracy of EKF-based CLS along different trajectories in simulations
Trajectory X Positioning Error, m Y Positioning Error, m Z Positioning Error, m
min max RMS min max RMS min max RMS
Case 1 (straight) 7.892e-4 0.8381 0.5769 0.01265 0.6452 0.3891 0.01802 1.109 0.942
Case 2 (zigzag) 0.1959 1.284 0.6168 0.1419 1.983 0.8118 0.02051 1.179 0.9636
Case 3 (ours) 0.004418 0.2121 0.07992 0.02308 0.2953 0.09616 0.02199 0.2339 0.05407
Table 4: Positioning Uncertainty of EKF-based CLS along different trajectories in simulations
[…]
Case 1 (straight) 104.789 62.469 121.943
Case 2 (zigzag) 74.959 73.587 135.399
Case 3 (ours) 61.466 58.792 27.316
[…]
During the simulation, executing the OPC consumes $0.5566\pm 0.1108s$ (mean $\pm$ std) per loop, requiring an average of 40 NLP solver iterations on average. This cost scales linearly with the number of stages $N$ in the horizon. Although this remains too slow for high-rate feedback control, it approaches acceptable upper margins for sparser online replanning.
6.2 Quadrotor Flight Experiments
Figure 6: A practical cooperative localization system based on UWB interrobot ranging mounted on a pair of F450 quadrotors, portrayed in flight during our flight test experiments
Table 5: Configuration of F450 Experimental Platforms
Type Name and Specifications
Airframe F450 (1.7 kg, 5“ props, 4500 mAh LiPo)
Autopilot Cube Orange (PX4 1.14.5)
Onboard Computer Jetson TX2 NX (JetPack 4.6)
GNSS unit (leader) UBlox M8N
UWB unit (follower) ESP32 UWB DWM3000
Next, we apply our OPC to formation flight experiments with quadrotors depicted in Fig. 6 and configured according to Table 5. The leader follows a straight trajectory, while the follower performs two test cases, excluding the zigzagging trajectory (Case 2 in simulations), as it contributed to worse positioning in simulations.
Case 1
A uniform and straight trajectory, planned ahead-of-time by a minimum-snap trajectory planner,
Case 2
An observability-aware trajectory, planned ahead-of-time by running the OPC as a sliding-window optimizer.
We use a range-only CLS interfaced with a UWB range system [^11] The C/C++ interface to the UWB sensor can be found at: https://github.com/FSC_Lab/dwm3000_ros to estimate the states of both quadrotors, so that the estimation performance can be compared between the two test cases.
Figures 7a and 7b show straight (Case 1) and observability-aware (Case 2) trajectories during flight tests. The observability-aware trajectory induces an orbiting motion in the follower, while the leader continues along a uniform straight path. This motion pattern yields significantly more excitation in the inter-vehicle range.
(a) Quadrotor trajectories in flight experiment Case 1.
(b) Quadrotor trajectories in flight experiment Case 2.
Figure 7: Trajectories executed by the quadrotor team in flight experiments Case 1 and 2 (ours)
[…]
Table 6: Positioning Uncertainty of EKF-based CLS along different trajectories in flight tests. Position estimation error, RMSE error, and uncertainty ($3\sigma$ envelope) are shown along each axis
[…]
Case 1 (straight) 119.0 122.6 153.5
Case 2 (ours) 53.97 49.61 59.81
In the absence of ground-truth data, the positioning error in Fig. 8 is relative to the postprocessed GNSS measurements, which have a positioning variance of approximately 0.3 m. However, this error is still broadly illustrative of the CLS performance. Our optimized trajectory in Case 2 maintained stable and bounded positioning error on both the X and the Y axes. In contrast, the straight trajectory in Case 1 exhibited asymmetric performance: better accuracy on the X-axis but significantly worse on the Y-axis, reflecting poor excitation in that dimension. This shows that the optimized trajectory guarantees an acceptable positioning performance evenly across all axes, which is consistent with our goal of maximizing the minimum improvement in precision along the least observable directions.
Together, our simulations and flight experiments demonstrate that our OPC can effectively improve the accuracy and precision of state estimation for the range-only quadrotor CLS. This improvement is achieved in the presence of realistic noise in simulations and practical model uncertainties in flight tests, even though our observability cost was derived assuming noiseless state evolution. This validates our approach and reinforces the utility of the OPC in applied settings.
(a) X-axis position error
(b) Estimated X-axis uncertainty ($3\sigma$ envelope)
(c) Y-axis position error
(d) Estimated Y-axis uncertainty ($3\sigma$ envelope)
(e) Z-axis position error
(f) Estimated Z-axis uncertainty ($3\sigma$ envelope)
Figure 8: Position estimation performance of EKF-based CLS during flight experiments Case 1 and 2 (ours); Squares and circles mark starting and end positions, respectively
7 Conclusion
We proposed an observability-aware controller to improve positioning precision for quadrotor drones that perform formation flight using range-only CL. This controller addresses the issue that a part of the system state is not directly observable through range measurements, and higher estimation precision can be achieved by a control strategy that maximizes observability. This controller builds on recent work on the LOG, which captures information about the system’s observability better than first-order quantities such as the FIM, CRLB, or the EKF covariance.
We derived STLOG, a novel point-wise approximation of the LOG. Using the STLOG, we formulate an observability-optimal control principle that maximizes the sum of the minimum improvements in estimation precision along a trajectory. We designed the OPC, a receding-horizon controller, based on this principle. The OPC refines the estimation in the least observable directions of the quadrotor state and improves the reliability of range-only CL.
We validated our approach using the accuracy and precision of the estimation as performance metrics when we applied the OPC to a range-only quadrotor CLS. Our Monte Carlo simulations showed consistent improvements in estimator performance along OPC-generated trajectories compared to both straight and artificially perturbed trajectories. These improvements were further supported by flight experiments, in which our OPC was deployed on a quadrotor team performing CL.
Appendix A Formulas regarding likelihood ratios of infinite-dimensional spaces
In Section 4, we required the computation of likelihood ratios of the form $p(\mathbf{y}(t)|\mathbf{x}_{0})/p(\mathbf{y}(t))$ (e.g. in Eq. (16)). These likelihood ratios can be formalized as Radon–Nikodym densities of the underlying probability measures on infinite-dimensional spaces. In this appendix we outline the construction of these measures using Gaussian measure theory in general. We assume familiarity with the measure-theoretic formulation of probability in this section, for which we refer to [25]. We also use established results in Gaussian measure theory extensively, for which we refer to [26, 27].
First, we state the main results of Gaussian measure theory on separable Hilbert spaces. Let $H$ be a separable Hilbert space with inner product $\langle\cdot,\cdot\rangle_{H}$. A Borel probability measure $\mu$ on $H$ is called a Gaussian measure if, for every continuous linear functional $l\in H^{*}$, the pushforward measure $l_{*}\mu$ on $\mathbb{R}$ is a one-dimensional Gaussian measure [^12] Here we define a one-dimensional Gaussian measure on $\mathbb{R}$ as a Borel measure $\mu$ whose Fourier transform is given as $\int_{\mathbb{R}}e^{-ikx}d\mu(x)=\exp(imk-qk^{2})$ for some real parameters $m\in\mathbb{R},q\in[0,+\infty)$, so it includes the Dirac measure in particular.. Moreover, we call the measure $\mu$ centered if $l_{*}\mu$ has zero mean for all $l\in H^{*}$.
The mean $m:H^{*}\rightarrow\mathbb{R}$ and covariance $\mathcal{K}:H^{*}\times H^{*}\rightarrow\mathbb{R}$ of the Gaussian measure $\mu$ is then defined as the continuous linear (resp. bilinear) functionals
$m(l)=\int_{H}l(y)\ d\mu(y),\quad\mathcal{K}(l,l^{\prime})=\int_{H}l(y-m)l^{\prime}(y-m)\ d\mu(y),$ (44)
for all $l,l^{\prime}\in H^{*}$. Here we identified $m\in H^{**}\cong H$ in Eq. (44). Moreover, the covariance $\mathcal{K}$ in Eq. (44) induces a bounded, self-adjoint and trace-class operator on $H$ through the canonical isomorphisms $H\cong H^{*}\cong H^{**}$, which we also denote as $\mathcal{K}$ by abuse of notation.
Conversely, let $\mathcal{K}:H\rightarrow H$ be a bounded, self-adjoint and trace-class operator. Then $\mathcal{K}$ defines a centered Gaussian measure $\mu_{0,\mathcal{K}}$ using the formulas in Eq. (44) with $m=0$. Consider the canonical map $\iota:H^{*}\rightarrow L^{2}(H,\mu_{0,\mathcal{K}})$, where we consider a linear functional of $H$ as a $\mu_{0,\mathcal{K}}$-square-integrable function. The closure of $\iota(H^{*})$ in $L^{2}(H,\mu_{0,\mathcal{K}})$ under the Hilbert space norm is known as the reproducing kernel Hilbert space (RKHS) $\mathcal{R}_{\mathcal{K}}$, and is itself a Hilbert space with the inner product
$\langle\iota(l),\iota(l^{\prime})\rangle=\int_{H}l(y)l^{\prime}(y)\ d\mu_{0,\mathcal{K}}(y)=\mathcal{K}(l,l^{\prime}).$ (45)
[…]
$l(h)=\int_{H}l(y)\hat{h}(y)\ d\mu_{0,\mathcal{K}}(y)\quad\text{for all $l\in H^{*}$},$ (46)
and we can show in fact that $h\mapsto\hat{h}$ is an isomorphism of Hilber spaces $\mathcal{H}_{\mathcal{K}}\cong\mathcal{R}_{\mathcal{K}}$ when the Cameron–Martin space is equipped with the following inner product:
$\langle h,k\rangle_{\mathcal{H}_{\mathcal{K}}}=\langle\hat{h},\hat{k}\rangle_{\mathcal{R}_{\mathcal{K}}}.$ (47)
The Cameron–Martin theorem characterizes when two Gaussian measures with the same covariance are absolutely continuous w.r.t. each other:
Theorem A.1 (Cameron–Martin).
The Gaussian measure $\mu_{h,\mathcal{K}}$ is absolutely continuous with respect to $\mu_{0,\mathcal{K}}$, if and only if $h\in\mathcal{H}_{\mathcal{K}}$. In this case, the Radon–Nikodym density is
$\frac{d\mu_{h,\mathcal{K}}}{d\mu_{0,\mathcal{K}}}(y)=\exp\left(\hat{h}(y)-\frac{1}{2}\langle h,h\rangle_{\mathcal{H}_{K}}\right),$ (48)
where $\hat{h}$ and $\langle h,h\rangle_{\mathcal{H}_{\mathcal{K}}}$ are given in Eqs. (46) and (47), respectively.
Now consider the spectral decomposition of $\mathcal{K}$: let ${\{e_{i}\}}_{i=1}^{\infty}$ be an orthonormal eigenbasis for $\mathcal{K}$ with corresponding eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$. If $\lambda_{i}>0$ for all $i$, we have the following expedient characterization:
  1. 1.
    $\mathcal{H}_{\mathcal{K}}$ is the set of elements $h$ satisfying $\sum_{i=1}^{\infty}\langle h,e_{i}\rangle^{2}/\lambda_{i}<\infty$;
  2. 2.
    The inner product in Eq. (47) is given explicitly as $\langle h,k\rangle_{\mathcal{H}_{\mathcal{K}}}=\langle\mathcal{K}^{-1/2}h,\mathcal{K}^{-1/2}k\rangle_{H}$;
  3. 3.
    The map $h\mapsto\hat{h}$ in Eq. (46) is given explicitly as $\hat{h}=\mathcal{K}^{-1}h$.
On the contrary, if $\mathcal{K}$ only has a finite number of non-zero eigenvalues, then $\mathcal{H}_{\mathcal{K}}=\text{range}(\mathcal{K})$ algebraically, and in particular we can only compute likelihood ratios when we constrain the mean in a finite-dimensional subspace.
In the following, we focus on the case where $\mathcal{K}$ is positive definite i.e. has no zero eigenvalues.
The Radon–Nikodym density in Eq. (48) is merely a $\mu_{0,\mathcal{K}}$-measurable function on $H$ for a fixed $h\in\mathcal{H}_{\mathcal{K}}$. An easy way to obtain more uniformity on both $h$ and $y\in H$ is to only consider mean displacements in a smaller subspace $\mathcal{H}_{\mathcal{K}^{2}}\subset\mathcal{H}_{\mathcal{K}}$. Since $\mathcal{H}_{\mathcal{K}^{2}}\hookrightarrow\mathcal{H}_{\mathcal{K}}\hookrightarrow H$ and $\mathcal{H}_{\mathcal{K}^{2}}\xrightarrow[]{\mathcal{\mathcal{K}}^{-1/2}}\mathcal{H}_{\mathcal{K}}\xrightarrow[]{\mathcal{K}^{-1/2}}H$ are continuous, we have:
Corollary A.1.1.
Let $\mathcal{K}$ be a positive definite trace-class operator on a separable Hilbert space $H$, and let $\mathcal{H}_{\mathcal{K}^{2}}$ be the Cameron–Martin space of $\mathcal{K}^{2}$. For $h\in\mathcal{H}_{\mathcal{K}^{2}}$, the Radon–Nikodym density of $\mu_{h,\mathcal{K}}$ with respect to $\mu_{0,\mathcal{K}}$ exists, and moreover is a continuous function on $\mathcal{H}_{K^{2}}\times H$ given by
$\frac{d\mu_{h,\mathcal{K}}}{d\mu_{0,\mathcal{K}}}(y)=F(h,y)=\exp\left(\langle y,\mathcal{K}^{-1}h\rangle_{H}-\frac{1}{2}\langle h,\mathcal{\mathcal{K}}^{-1}h\rangle_{H}\right).$ (49)
Now consider a finite-dimensional space $X\subset\mathbb{R}^{n}$ and a continuous function $\mathcal{A}:X\rightarrow\mathcal{H}_{\mathcal{K}^{2}}$, and also let $\nu$ be a Borel probability measure on $X$. The space $X$ should be considered as a finite-dimensional parametrization of possible means for the Gaussian measure on $H$, and $\nu$ our uncertainty of this parameter. We define the joint probability distribution on $X\times H$ to be the measure
$\gamma(x,y)=F(\mathcal{A}(x),y)(\nu\otimes\mu_{0,\mathcal{K}})(x,y),$ (50)
where $F$ is the positive function given by Eq. (49) and $\otimes$ denotes the product measure. Since $(x,y)\mapsto F(\mathcal{A}(x),y)$ is continuous, it is Borel measurable and $\nu\otimes\mu_{0,\mathcal{K}}$-integrable. We check by Tonelli’s theorem that
$\int_{X\times H}F(\mathcal{A}(x),y)\ d(\nu\otimes\mu_{0,\mathcal{K}})(x,y)=$ $\int_{X}\left(\int_{H}F(\mathcal{A}(x),y)\ d\mu_{0,\mathcal{K}}(y)\right)d\nu(x)=\int_{X}\left(\int_{H}d\mu_{\mathcal{A}(x),\mathcal{K}}(y)\right)d\nu(x)$
$=$ $\int_{X}1\ d\nu=1,$
Using the Cameron–Martin formula in Eq. (48) in the penultimate step, we can verify that the conditional measure $\gamma|_{x}$ of $\gamma$ in Eq. (50) at a given $x\in X$ and the marginal measure $\bar{\gamma}$, both measures on $H$, are given respectively by
$\gamma|_{x}(y)=F(\mathcal{A}(x),y)\mu_{0,\mathcal{K}}(y)=\mu_{\mathcal{A}(x),\mathcal{K}}(y),\quad\bar{\gamma}(y)=\left(\int_{X}F(\mathcal{A}(x),y)\ d\nu(x)\right)\mu_{0,K}(y),$ (51)
from which it follows that both $\gamma|_{x}$ and $\bar{\gamma}$ have positive Radon–Nikodym densities with respect to $\mu_{0,\mathcal{K}}$, and therefore are absolutely continuous with respect to each other. We summarize our findings in Proposition A.2:
Proposition A.2.
Consider a finite-dimensional parameter space $X\subset\mathbb{R}^{n}$ and a Borel measure $\nu$ on $X$. Let $\mu_{\mathcal{A}(x),\mathcal{K}}$ be a family of Gaussian measures on a Hilbert space $H$, with the same positive-definite trace-class covariance $\mathcal{K}$, and mean parametrized by a continuous map $\mathcal{A}:X\rightarrow\mathcal{H}_{\mathcal{K}^{2}}$. Consider the following two distributions on $H$:
  1. 1.
    $\gamma|_{x}=\mu_{\mathcal{A}(x),\mathcal{K}}$, the conditional distribution of $y\in H$ for a given $x\in X$,
  2. 2.
    $\bar{\gamma}$, the marginal distribution of $y\in H$ obtained from the marginalizing the joint distribution on $X\times H$,
then $\gamma|_{x}$ and $\bar{\gamma}$ are absolutely continuous with respect to each other, with Radon–Nikodym density
$\frac{d\gamma|_{x}}{d\bar{\gamma}}(y)$ $=C(y)F(\mathcal{A}(x),y),\quad C(y)=\left(\int_{X}F(\mathcal{A}(x),y)\ d\nu(x)\right)^{-1},$ (52)
where $F$ is given by Eq. (49) and $C(y)$ is independent of $x\in X$.
A useful special case is when $H=L^{2}([0,T],\mathbb{R}^{p})$, the Hilbert space of square-intergrable $\mathbb{R}^{p}$-valued functions on the closed interval $[0,T]$. Our convention for the Hilbert space inner product is
$\left\langle\mathbf{f}(t),\mathbf{g}(t)\right\rangle_{L^{2}([0,T],\mathbb{R}^{p})}=\frac{1}{T}\int_{0}^{T}\mathbf{f}^{\top}(t)\mathbf{g}(t)\ dt.$ (53)
The continuity of the map $\mathcal{A}:X\rightarrow\mathcal{H}_{\mathcal{K}^{2}}$ required by Proposition A.2 is often more expediently checked by introducing some intermediate Sobolev space $H^{s}([0,T],\mathbb{R}^{p})$.
Corollary A.2.1.
Consider a finite-dimensional parameter space $X\subset\mathbb{R}^{n}$ and a Borel measure $\nu$ on $X$. Let $\mathcal{K}$ be a positive-definite trace-class operator in the Hilbert space $L^{2}([0,T],\mathbb{R}^{p})$, with unbounded inverse $\mathcal{K}^{-1}$. Suppose that there exists a Sobolev space $H^{s}([0,T],\mathbb{R}^{p})$ for some $s\in\mathbb{Z}^{+}$ such that:
  1. 1.
    The map $\mathcal{A}:X\rightarrow H^{s}([0,T],\mathbb{R}^{p})$ is continuous;
  2. 2.
    There exists a constant $C>0$ such that
    $\norm{\mathcal{K}^{-1}\mathbf{h}}_{L^{2}([0,T],\mathbb{R}^{p})}\leq C\norm{\mathbf{h}}_{H^{s}([0,T],\mathbb{R}^{p})}\quad\text{for all $\mathbf{h}\in H^{s}([0,T],\mathbb{R}^{p})$}$ (54)
    where $\norm{\cdot}_{H^{s}([0,T],\mathbb{R}^{p})}$ denotes the corresponding Sobolev norm;
Then for all $x\in X$, the measures
  1. 1.
    $\gamma|_{x}$, the Gaussian distribution on $L^{2}([0,T],\mathbb{R}^{p})$ with mean $\mathcal{A}(x)$ and covariance $\mathcal{K}$, and
  2. 2.
    $\bar{\gamma}$, the marginal distribution on $L^{2}([0,T],\mathbb{R}^{p})$ obtained from marginializing $x$ in the joint distribution in Eq. (50),
are absolutely continuous with respect to each other, with Radon–Nikodym density at $\mathbf{y}(t)$
$\frac{d\gamma|_{x}}{d\bar{\gamma}}(\mathbf{y}(t))$ $=C(\mathbf{y}(t))\exp\left(\frac{1}{T}\int_{0}^{T}\left(\mathbf{y}^{\top}(t)\left(\mathcal{K}^{-1}\mathcal{A}(x)\right)(t)-\frac{1}{2}\mathcal{A}(x)^{\top}(t)\left(\mathcal{K}^{-1}\mathcal{A}(x)\right)(t)\right)\ dt\right),$ (55)
where $C(\mathbf{y}(t))$ is a constant independent of $x$.
Proof.
Simply note that the condition in Eq. (54) is equivalent to the continuity of the embedding $\iota:H^{s}([0,T],\mathbb{R}^{p})\hookrightarrow\mathcal{H}_{\mathcal{K}^{2}}$. Then $\mathcal{A}:X\rightarrow H^{s}([0,T],\mathbb{R}^{p})\hookrightarrow\mathcal{H}_{\mathcal{K}^{2}}$ is continuous, and the assumptions of Proposition A.2 are met. ∎
Remark.
Condition 1 of Corollary A.2.1 is often the ‘best-case’ result we can hope for if $\mathcal{A}$ is a map that involves the solution of ordinary differential equations. In practice, we can assume that $\mathcal{A}:X\rightarrow C^{\infty}([0,T],\mathbb{R}^{p})$ is continuous with respect to the Fréchet topology in the image, so in particular the map is continuous into Sobolev spaces of all orders.
Example A.1.
We give a recipe for constructing covariance operators $\mathcal{K}$ satisfying condition 2 of Corollary A.2.1. We discuss only the case for scalar-valued functions, i.e. $p=1$. The vector-valued case follows easily.
Let ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ be the orthonormal Legendre polynomials on $[0,T]$ under the inner product in Eq. (53). For a given function $h\in L^{2}([0,T])$, let $h_{n}=\langle h,e_{n}\rangle_{L^{2}([0,T])}$ be its $n$-th Fourier–Legendre coefficient. From [28], inequality (5.4.12), it follows that
$h\in H^{s}([0,T])\Rightarrow\lvert h_{n}\rvert\leq An^{-s}\norm{h}_{H^{s}([0,T])},$ (56)
where $A$ is a constant independent of $h$ and $n$. In particular, if $h(t)$ is smooth, then $h_{n}={O}(n^{-s})$ for all $s>0$.
Now consider the operator $\mathcal{K}$, defined spectrally as the operator with eigenbasis ${\left\{e_{n}(t)\right\}}_{n=0}^{\infty}$ and eigenvalues ${\left\{\lambda_{n}\right\}}_{n=0}^{\infty}$, so that
  1. 1.
    $\lambda_{n}>0$ for all $n$,
  2. 2.
[…]
  3. 3.
    For some $a>0$, $\alpha>1$ and $N>0$, we have, for all $n>N$, that $\lambda_{n}\geq an^{-\alpha}$,
Then $\mathcal{K}$ is a valid covariance operator, and moreover is positive definite. We also have:
$\norm{\mathcal{K}^{-1}h}_{L^{2}([0,T])}\leq\max(\lambda_{0},\ldots,\lambda_{N})\norm{h}_{L^{2}([0,T])}+\frac{1}{a}\sum_{n=N+1}^{\infty}n^{\alpha}\lvert h_{n}\rvert$, (57)
whenever $\mathcal{K}^{-1}h\in L^{2}([0,T])$. Comparing Eqs. (56) and (57) shows that, for any integer $s>\alpha+1$, the condition $h\in\norm{h}_{H^{s}([0,T])}$ is sufficient for the right-hand side of Eq. (57) to converge, and moreover in such a case an inequality of the form Eq. (54) holds.
Remark.
An analogous construction to Example A.1 also works for the Fourier basis instead of the Legendre basis on $L^{2}([0,T])$, but one has to restrict to periodic functions when considering the Sobolev and Fréchet spaces on $[0,T]$.
Appendix B Asymptotic size of the minimum eigenvalue of the STLOG
We derive Proposition 5.2 in this appendix. This asymptotic result relates the STLOG in Eq. (34) to the local observability index $r_{*}$ (14) defined as a bridge to rank-testing observability in Sec. 3. We consider a fixed operating point of the system $(\mathbf{x},\mathbf{u})$, and adopt several inessential simplifications. Specifically, we set $\boldsymbol{\Sigma}=\mathbf{1}$, and simply denote the order-$r$ STLOG from Eq. (34) by $\mathbf{W}^{(r)}$ without the subscript $\boldsymbol{\Sigma}^{-1}$.
Theorem B.1.
Let $(\mathbf{x},\mathbf{u})$ be given and $r_{*}(\mathbf{x},\mathbf{u})$ be its local observability index. Consider the minimum eigenvalue of the order-$r$ STLOG $\mathbf{W}^{(r)}(\mathbf{x},\mathbf{u};\Delta T)$.
  1. 1.
    If $r<r_{*}(\mathbf{x},\mathbf{u})$, then $\lambda_{\min}(\mathbf{W}^{(r)}(\mathbf{x},\mathbf{u};\Delta T))=0$ .
  2. 2.
    If $r\geq r_{*}(\mathbf{x},\mathbf{u})$, then for all sufficiently small $\Delta T$, there exists positive constants $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ independent of $\Delta T$ such that
    $C_{1}(\mathbf{x},\mathbf{u})\Delta T^{2r_{*}(\mathbf{x},\mathbf{u})+1}\leq\lambda_{\min}(\mathbf{W}^{(r)}(\mathbf{x},\mathbf{u};\Delta T))\leq C_{2}(\mathbf{x},\mathbf{u})\Delta T^{2r_{*}(\mathbf{x},\mathbf{u})+1}.$ (58)
    In particular, there exists some $C\in[C_{1},C_{2}]$ and a sequence $\Delta T_{k}\to 0$ such that $\lim_{k\to\infty}\frac{\lambda_{\min}(\mathbf{W}^{(r)}(\mathbf{x},\mathbf{u};\Delta T_{k}))}{\Delta T_{k}^{2r_{*}(\mathbf{x},\mathbf{u})+1}}=C$.
Proof.
We prove Part 1 in Proposition B.3, and provide expressions for $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ in Propositions B.4 and B.7 respectively. ∎
In Section 3, we reviewed that the system is locally weakly observable at order $r$ if $\text{rank }\boldsymbol{\mathcal{O}}^{(r)}=\dim\mathcal{X}$. We introduce the following subspace:
$\mathcal{H}_{k}=\bigcap_{i=0}^{k}\ker\left(DL_{\mathbf{f}}^{i}\mathbf{h}\right)=\ker\boldsymbol{\mathcal{O}}^{(k)}.$ (59)
Clearly, $\mathcal{H}_{k-1}\supset\mathcal{H}_{k}$. This decreasing sequence of subspaces reaches $\{0\}$ precisely at the local observability index $r_{*}$. We collect some elementary consequences of this definition into the following Lemma:
Lemma B.2.
  1. 1.
    $\mathcal{H}_{r}=\{0\}$ $\Leftrightarrow$ the system is locally observable at order $r$.
  2. 2.
    The local observability index $r_{*}(\mathbf{x},\mathbf{u})$ is the smallest value of $r$ such that $\mathcal{H}_{r}=\{0\}$.
  3. 3.
    In particular, if $r_{*}>0$, then $\mathcal{H}_{r_{*}-1}$ is nonzero.
We rewrite the STLOG from Eq. (34) in terms of $\boldsymbol{\mathcal{O}}^{(r)}$ before evaluating the integral (27). For any $\mathbf{v},\tilde{\mathbf{v}}\in\mathcal{X}$, [^13] In Eq. (60), if instead of the standard inner product on $\mathcal{Y}$ we use the inner product induced by $\boldsymbol{\Sigma}^{-1}$, we recover the expressions in Eqs. (27) and (34).
$\mathbf{v}^{\top}\mathbf{W}^{(r)}\tilde{\mathbf{v}}=\int_{0}^{\Delta T}$ ${\left(\begin{bmatrix}\mathbf{1}&t\mathbf{1}&\cdots&\frac{t^{r}}{r!}\mathbf{1}\end{bmatrix}{\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}\right)}^{\top}\left(\begin{bmatrix}\mathbf{1}&t\mathbf{1}&\cdots&\frac{t^{r}}{r!}\mathbf{1}\end{bmatrix}{\boldsymbol{\mathcal{O}}^{(r)}}\tilde{\mathbf{v}}\right)\ dt,$ (60)
making $\mathbf{W}^{(r)}$ positive semi-definite by construction. This is linked to local observability by Proposition B.3.
Proposition B.3.
The system is not locally observable at order $r$ if and only if $\lambda_{\min}(\mathbf{W}^{(r)})=0$.
Proof.
Suppose the system is not locally observable at order $r$. By Lemma B.2, there exists some nonzero $\mathbf{v}\in\mathcal{H}_{r}$. In particular, ${\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}=0$. Substituting such $\mathbf{v}$ into Eq. (60) gives $\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}=0$, so $\lambda_{\min}(\mathbf{W}^{(r)})=0$.
Conversely, if $\lambda_{\min}(\mathbf{W}^{(r)})=0$, then there exists some nonzero $\mathbf{v}\in\mathcal{X}$ such that $\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}=0$. Rewriting Eq. (60):
$\int_{0}^{\Delta T}\norm{\begin{bmatrix}\mathbf{1}&t\mathbf{1}&\cdots&\frac{t^{r}}{r!}\mathbf{1}\end{bmatrix}{\boldsymbol{\mathcal{O}}^{(r)}}{\mathbf{v}}}^{2}\ dt=0,$ (61)
from which we deduce $\begin{bmatrix}\mathbf{1}&t\mathbf{1}&\cdots&\frac{t^{r}}{r!}\mathbf{1}\end{bmatrix}{\boldsymbol{\mathcal{O}}^{(r)}}{\mathbf{v}}=0$ for all $t\in[0,\Delta T]$, which implies ${\boldsymbol{\mathcal{O}}^{(r)}}{\mathbf{v}}=0$, thus we conclude the system is not locally observable at order $r$. ∎
Next, we derive the upper and lower bounds in Eq. (B.1) in Theorem B.1. Let $r_{*}$ be the local observability index. We only consider $\mathbf{W}^{(r)}$ with $r\geq r_{*}$ in the following, and assume $r_{*}>0$ without loss of generality.
We first derive the upper bound in Eq. (B.1). We leverage the identity on analytical minima of a quadratic form over a unit sphere $\lambda_{\min}(\mathbf{W}^{(r)})=\min_{\mathbf{v}\in\mathcal{X},\ \norm{\mathbf{v}}=1}\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}$. We thus seek an appropriate unit vector $\mathbf{v}$ that demonstrates the upper bound in Eq. (B.1). Any unit vector $\mathbf{v}\in\mathcal{H}_{r^{*}-1}$ satisfies
$\left(DL_{\mathbf{f}}^{r_{*}}\mathbf{h}\right)\mathbf{v}\neq 0,\ \left(DL_{\mathbf{f}}^{k}\mathbf{h}\right)\mathbf{v}=0\ \text{for all }k<r_{*}.$ (62)
In particular, $\norm{(DL_{\mathbf{f}}^{r_{*}}\mathbf{h})\mathbf{v}}=\norm{{\boldsymbol{\mathcal{O}}^{(r_{*})}}{\mathbf{v}}}$. We pick a unit vector $\mathbf{v}$ that minimizes $\norm{(DL_{\mathbf{f}}^{r_{*}}\mathbf{h})\mathbf{v}}$ and obtain:
$\norm{(DL_{\mathbf{f}}^{r_{*}}\mathbf{h})\mathbf{v}}=\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r^{*}-1}}\right)>0,$ (63)
where $|_{\mathcal{H}_{r^{*}-1}}$ denotes the restriction to the subspace $\mathcal{H}_{r^{*}-1}$. Now we evaluate $\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}$ using our selected $\mathbf{v}$:
$\begin{bmatrix}\mathbf{1}&t\mathbf{1}&\cdots&\frac{t^{r}}{r!}\mathbf{1}\end{bmatrix}{\boldsymbol{\mathcal{O}}^{(r)}}{\mathbf{v}}=\frac{t^{r_{*}}}{r_{*}!}\left(DL_{\mathbf{f}}^{r_{*}}\mathbf{h}\right)\mathbf{v}+t^{r_{*}+1}\mathbf{a}(\mathbf{v},t),$ (64)
for some polynomial $\mathbf{a}(\mathbf{v},t)$. Substituting Eq. (64) into Eq. (60) and evaluating the integral gives
$\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}=\frac{\Delta T^{2r_{*}+1}\norm{\left(DL_{\mathbf{f}}^{r_{*}}\mathbf{h}\right)\mathbf{v}}^{2}}{{(r_{*}!)}^{2}(2r_{*}+1)}+\Delta T^{2r_{*}+2}\alpha(\mathbf{v},\Delta T),$ (65)
[…]
$\lambda_{\min}(\mathbf{W}^{(r)})\leq\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}\leq\Delta T^{2r_{*}+1}\left(B+A\Delta T\right).$ (66)
For any $\epsilon>0$, choosing $\Delta T<\epsilon B/A$ gives:
Proposition B.4.
For all $\epsilon>0$, there exists $\delta>0$ such that
$\lambda_{\min}(\mathbf{W}^{(r)})\leq\frac{(1+\epsilon)\Delta T^{2r_{*}+1}\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r^{*}-1}}\right)}{{(r_{*}!)}^{2}(2r_{*}+1)},$ (67)
for all $\Delta T<\delta$.
Next we derive the lower bound in (58). It is convenient to rewrite the STLOG (34) into block matrix form
[…]
where $\boldsymbol{\Lambda}^{(r)}$ is the block diagonal matrix
$\boldsymbol{\Lambda}^{(r)}=\operatorname*{diag}\begin{pmatrix}\mathbf{1}&\Delta T\mathbf{1}&\cdots&\frac{\Delta T^{r}}{r!}\mathbf{1}\end{pmatrix}$, (69)
[…]
Lemma B.5.
Let $\mathbf{v}\in\mathcal{X}$ be a unit vector. Then for $\Delta T\leq 1$,
$\norm{\boldsymbol{\Lambda}^{(k)}{\boldsymbol{\mathcal{O}}^{(k)}}\mathbf{v}}\geq\frac{\Delta T^{k}}{k!}\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(k)}}\right).$ (71)
Proof.
By substituting the value of $\sigma_{\min}\left(\boldsymbol{\Lambda}^{(r)}\right)$ from (69) and $\norm{\mathbf{v}}=1$ in the last step, we get
$\norm{\boldsymbol{\Lambda}^{(k)}{\boldsymbol{\mathcal{O}}^{(k)}}\mathbf{v}}\geq\sigma_{\min}\left(\boldsymbol{\Lambda}^{(k)}{\boldsymbol{\mathcal{O}}^{(k)}}\right)\norm{\mathbf{v}}\geq\sigma_{\min}\left(\boldsymbol{\Lambda}^{(k)}\right)\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(k)}}\right)\norm{\mathbf{v}}=\frac{\Delta T^{k}}{k!}\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(k)}}\right),$
The bound (71) admits a straightforward but important improvement for $r\geq r_{*}$:
Lemma B.6.
Let $\mathbf{v}\in\mathcal{X}$ be a unit vector, and let $r\geq r_{*}$. Then for $\Delta T\leq 1$,
$\norm{\boldsymbol{\Lambda}^{(r)}{\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}}\geq\frac{\Delta T^{r_{*}}}{r_{*}!}\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right),$ (72)
where $\sigma_{\min}({\boldsymbol{\mathcal{O}}^{(r_{*})}})$ is positive and independent of $\Delta T$.
Proof.
Write the vector $\boldsymbol{\Lambda}^{(r)}{\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}$ as a block of two vectors
$\boldsymbol{\Lambda}^{(r)}{\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}=\begin{bmatrix}\boldsymbol{\Lambda}^{(r_{*})}{\boldsymbol{\mathcal{O}}^{(r_{*})}}\mathbf{v}\\ \mathbf{w}\end{bmatrix}$,
where $\mathbf{w}$ is a $(r-r_{*})\dim\mathcal{Y}\times 1$ column vector. Now
$\norm{\boldsymbol{\Lambda}^{(r)}{\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}}=\sqrt{\norm{\boldsymbol{\Lambda}^{(r_{*})}{\boldsymbol{\mathcal{O}}^{(r_{*})}}\mathbf{v}}^{2}+\norm{\mathbf{w}}^{2}}\geq\norm{\boldsymbol{\Lambda}^{(r_{*})}{\boldsymbol{\mathcal{O}}^{(r_{*})}}\mathbf{v}}\geq\frac{\Delta T^{r_{*}}}{r_{*}!}\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right),$
where we used the bound (71) for $k=r_{*}$ in the last step. ∎
Proposition B.7.
For all $\Delta T\leq 1$, we have
$\lambda_{\min}(\mathbf{W}^{(r)})\geq\frac{a^{(r)}\Delta T^{2r_{*}+1}\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right)}{{(r_{*}!)}^{2}},$ (73)
where $a^{(r)}>0$ is the minimum eigenvalue of the Hilbert matrix of order $r$ (70).
Proof.
[…]
$\mathbf{v}^{\top}\mathbf{W}^{(r)}\mathbf{v}\geq\Delta Ta^{(r)}\norm{\boldsymbol{\Lambda}^{(r)}{\boldsymbol{\mathcal{O}}^{(r)}}\mathbf{v}}^{2}\geq\Delta Ta^{(r)}{\left(\frac{\Delta T^{r_{*}}}{r_{*}!}\sigma_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right)\right)}^{2},$
using (72) in the last step. Simplifying and minimizing over unit vectors $\mathbf{v}\in\mathcal{X}$ gives the inequality (73). ∎
Combining Propositions B.4B.7, we have, for all sufficiently small $\Delta T$,
$\alpha\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}\right)\leq\frac{\lambda_{\min}(\mathbf{W}^{(r)})}{\Delta T^{2r_{*}+1}}\leq\beta\sigma^{2}_{\min}\left({\boldsymbol{\mathcal{O}}^{(r_{*})}}|_{\mathcal{H}_{r^{*}-1}}\right),$ (74)
where $\alpha,\beta$ are constants. Thus, we proved the inequality (58) in Theorem B.1 by explicitly constructing $C_{1}(\mathbf{x},\mathbf{u}),C_{2}(\mathbf{x},\mathbf{u})$ in terms of $\boldsymbol{\mathcal{O}}^{(r_{*})}$ only. The last part of Theorem B.1 follows from the Bolzano—Weierstrass theorem.
Appendix C Choice of STLOG order for quadrotor CLS
Proposition C.1.
Consider a dynamical system in Eqs. (2) and (3), whose state space can be partitioned into $\mathcal{X}=\mathcal{Z}\times\mathcal{Q}$, with state $\mathbf{x}=(\mathbf{z},\mathbf{q})$, $\mathbf{z}\in\mathcal{Z}$, $\mathbf{q}\in\mathcal{Q}$, such that the dynamics and the observations can be decomposed into:
$\begin{bmatrix}\dot{\mathbf{z}}\\ \dot{\mathbf{q}}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{f}_{z}(\mathbf{z},\mathbf{q},\mathbf{u})\\ \mathbf{f}_{q}(\mathbf{q},\mathbf{u})\end{bmatrix},\quad\mathbf{y}=\mathbf{h}(\mathbf{z})=\begin{bmatrix}\mathbf{d}(\mathbf{z},\mathbf{q})\\ \mathbf{a}(\mathbf{q})\end{bmatrix}$, (75)
where $\mathbf{f}_{z}$ and $\mathbf{f}_{q}$ are components of the dynamics $\mathbf{f}=(\mathbf{f}_{z},\mathbf{f}_{q})$, and $\mathbf{d}:\mathcal{Z}\times\mathcal{Q}\rightarrow\mathbb{R}^{s}$ and $\mathbf{a}:\mathcal{Q}\rightarrow\mathbb{R}^{p-s}$ are components for the observation map $\mathbf{h}:\mathcal{X}\rightarrow\mathcal{Y}=\mathbb{R}^{p}$. For all $(\mathbf{x},\mathbf{u})$, the local observability index $r_{*}$ in Eq. (14) for the system in Eq. (75) satisfies $s(r_{*}(\mathbf{x},\mathbf{u})+1)\geq\dim(\mathcal{Z})$ .
Proof.
The iterated Lie derivatives of $\mathbf{h}$ along $\mathbf{f}$ can be evaluated inductively as
$L_{\mathbf{f}}\mathbf{h}(\mathbf{z},\mathbf{q})=L_{\mathbf{f}}(\mathbf{d}(\mathbf{z},\mathbf{q}),\mathbf{a}(\mathbf{q}))=\left(L_{\mathbf{f}_{z}}+L_{\mathbf{f}_{q}}\right)\begin{bmatrix}\mathbf{d}(\mathbf{z},\mathbf{q})\\ \mathbf{a}(\mathbf{q})\end{bmatrix}=\begin{bmatrix}\mathbf{d}^{(1)}(\mathbf{z},\mathbf{q})\\ \mathbf{a}^{(1)}(\mathbf{q})\end{bmatrix},\quad L^{i}_{\mathbf{f}}\mathbf{h}(\mathbf{z},\mathbf{q})=\begin{bmatrix}\mathbf{d}^{(i)}(\mathbf{z},\mathbf{q})\\ \mathbf{a}^{(i)}(\mathbf{q})\end{bmatrix}$, (76)
for some functions $\mathbf{d}^{(i)}$ and $\mathbf{a}^{(i)}$, where $\mathbf{a}^{(i)}$ has no $\mathbf{z}$-dependence. We next evaluate the gradients of the Lie derivatives:
$DL^{i}_{\mathbf{f}}\mathbf{h}=\begin{bmatrix}D_{\mathbf{z}}\mathbf{d}^{(i)}&D_{\mathbf{q}}\mathbf{d}^{(i)}\\ D_{\mathbf{z}}\mathbf{a}^{(i)}&D_{\mathbf{q}}\mathbf{a}^{(i)}\\ \end{bmatrix}=\begin{bmatrix}D_{\mathbf{z}}\mathbf{d}^{(i)}&D_{\mathbf{q}}\mathbf{d}^{(i)}\\ \mathbf{0}&D_{\mathbf{q}}\mathbf{a}^{(i)}\\ \end{bmatrix}$. (77)
Finally, we build the observability matrix:
$\boldsymbol{\mathcal{O}}^{(r)}=\begin{bmatrix}D_{\mathbf{z}}\mathbf{d}&D_{\mathbf{q}}\mathbf{d}\\ \mathbf{0}&D_{\mathbf{q}}\mathbf{a}\\ \vdots&\vdots\\ D_{\mathbf{z}}\mathbf{d}^{(r)}&D_{\mathbf{q}}d^{(r)}\\ \mathbf{0}&D_{\mathbf{q}}\mathbf{a}^{(r)}\\ \end{bmatrix}$. (78)
The submatrix of $\boldsymbol{\mathcal{O}}^{(r)}$ consisting the first $\dim\mathcal{Z}$ columns has no more than $s(r+1)$ nonzero rows, so the row span of $\boldsymbol{\mathcal{O}}^{(r)}$ will always be deficient if $s(r+1)<\dim\mathcal{Z}$.
Corollary C.1.1.
For the quadrotor CLS system in Eqs. (4) and (7), $r_{*}\geq 5$.
Proof.
Our quadrotor CLS system admits the decomposition in Proposition C.1, with:
  1. 1.
    $\mathcal{Z}$ is the space for the $6$-dimensional translational state including position and velocity $\mathbf{z}=(\mathbf{r},\mathbf{v})$;
  2. 2.
    $\mathcal{Q}$ is the space for the rotational state $\mathbf{q}$;
  3. 3.
    $\mathbf{d}$ is the $1$-dimensional range-only measurement $(s=1)$;
  4. 4.
    $\mathbf{a}(\mathbf{q})=\mathbf{q}$ is the direct measurement of the rotational state;
So applying Proposition C.1 gives $1\times(r_{*}+1)\geq\dim\mathcal{Z}=6$, as required. ∎
Acknowledgements
We would like to acknowledge the sponsorship by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the Discovery Grant RGPIN-2023-05148. We also acknowledge our summer student, Tommy Zhang, for contributing to flight testing and hardware interfacing of quadrotor platforms.